### v2

# Low-Cost Architecture and Efficient Pattern Synthesis for Polarimetric Phased Array Based on Polarization Coding Reconfigurable Elements

Abstract
Polarimetric phased arrays (PPAs) enhance radar target detection and anti-jamming capabilities. However, the dual transmit/receive (T/R) channel requirement leads to high costs and system complexity. To address this, this paper introduces a polarization-coding reconfigurable phased array (PCRPA) and associated pattern synthesis techniques to reduce PPA costs while minimizing performance degradation. Each PCRPA element connects to a single T/R channel and incorporates two-level RF switches for real-time control of polarization states and waveforms. By adjusting element codes and excitation weights, the PCRPA can generate arbitrarily polarized and dual-polarized beams. Efficient beam pattern synthesis methods are also proposed, featuring novel optimization constraints derived from theoretical and analytical analysis of PCRPAs. Simulations demonstrate that the approach achieves low cross-polarization and sidelobe levels comparable to conventional architectures within the scan range, particularly for large arrays. However, the channel reduction inevitably incurs power and directivity loss. Experiments conducted on an $8\times 8$ X-band array antenna validate the effectiveness of the proposed system. The PCRPA and synthesis methods are well-suited for large-scale PPA systems, offering significant cost-effectiveness while maintaining good sidelobe suppression and polarization control performance.
Index Terms:
Polarimetric phased array radar, low-cost phased array, polarization pattern synthesis, cross-polarization level (XPL).
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
Polarimetric phased arrays (PPA) tend to be adopted by many radar systems to enhance their measurement performance within complex environments. The extra dimension of polarization can provide detailed information to detect targets in the presence of strong clutter[1, 2, 3, 4, 5], interference[6, 7, 8, 9, 10], or other adjacent objects[11, 12, 13]. These applications rely on the beamforming capabilities of dual- or arbitrarily polarized patterns, which require independent channels of orthogonal polarization states, usually horizontal (H) and vertical (V). Therefore, conventional PPA architectures generally consist of dual-polarized antenna elements with transmit/receive (T/R) channels connected to each of their ports[14, 15, 16, 17]. Compared with single-polarized phased arrays of the same size, the number of channels in the T/R modules of PPAs is doubled.
The conventional architectures are high-configuration but add to the burden of manufacturing the PPA systems, especially for large-scale arrays. A typical cost distribution for phased arrays given by [18] shows that the T/R modules can account for half the system cost, whose components grow proportionally with the number of elements. Meanwhile, many recently reported advanced PPA radar systems are on the scale of 1000-element or even bigger[19, 20, 21] to obtain high space resolution and more degrees of freedom (DOFs) for signal processing. Therefore, for these large-scale PPA systems, the heavy use of T/R channels under conventional architectures will make the total cost and complexity fairly high. And this also reveals the practical opportunity for cost reduction.
Over the past few decades, techniques like sparse arrays, thinned arrays, and subarrays have been thoroughly studied. They can significantly reduce the cost and complexity of phased arrays and directly apply to PPAs[22, 23]. However, these array designs are relatively fixed and static, so the configuration must be determined before manufacturing and cannot be changed. This will limit some crucial performance of radar systems, especially the power and scan range. In addition, polarization reconfigurable antennas are generally designed with single ports, requiring half the T/R channels compared to dual-polarized antennas[24, 25, 26]. If they are employed as PPA elements, the system cost and complexity will also be significantly reduced. Novel algorithms considering element polarization as variables can be proposed to ensure the performance of polarization pattern synthesis during the working process. Since the polarization states are adjusted in real-time, the design can be regarded as a dynamic thinned array with polarization as a new DOF.
Current research related to this issue can be divided into two schemes. The first has thoroughly investigated the application of polarization-reconfigurable elements in metasurfaces[27, 28], transmitarrays[29], and other non-electronically scanned array antennas[30, 31]. These works are instructive, but their array architectures and pattern synthesis models differ from the PPAs. The second scheme explores the idea of mixed polarized elements in PPA pattern synthesis but does not involve specific architecture design or beamforming requirements[32, 33, 34, 35, 36]. For example, in [33, 34], a signal processing method called cross-polarization canceller (XPC) uses a selection of cross-polarized elements to improve the cross-polarization level (XPL) of beam patterns by canceling the corresponding field components. However, it is based on the conventional architecture and requires the phased array to be fully digital. Several architecture models have been reported in [35, 36], but they do not cover the commonly used modes of PPA radars, such as alternately transmitting and simultaneously receiving (ATSR) and simultaneously transmitting and simultaneously receiving (STSR).
This paper proposes an architecture and corresponding polarization pattern synthesis methods for PPA radars to reduce the cost and complexity while ensuring beamforming performance. The proposed architecture employs elements with polarization reconfigurable ability. Given that the polarization states of each element are limited and independently controllable, this architecture is termed the polarization coding reconfigurable phased array (PCRPA). In the hardware design, each T/R channel has two switches: the antenna-connected switch for the antenna polarization state selection and the transceiver-connected switch for the waveform selection. This paper proposes novel constraints, derived from antenna radiation and array geometry analysis, to optimize pattern performance and enhance algorithm efficiency. Combined with classical optimization algorithms, efficient synthesis methods for the polarization pattern of the PCRPA are proposed. The main contributions of this paper are listed below.
  1. 1.
    A novel array architecture considering multiple applications for PPA radars is proposed. The designed architecture enables the real-time coding of polarization states and waveforms for reconfigurable antenna elements, allowing for the synthesis of arbitrarily polarized and dual-polarized beam patterns. Compared to conventional architectures, the proposed design offers a cost-effective advantage by reducing the number of active channels in the T/R modules by half.
  2. 2.
    The pattern performance of arbitrarily polarized and dual-polarized beams is analyzed through simulations. Two pattern synthesis methods are proposed based on the different performance to efficiently control power and polarization properties, where elements’ polarization states are introduced as variables.
  3. 3.
    The effectiveness of the proposed architecture and pattern synthesis methods is verified through electromagnetic simulations and experiments. A detailed discussion is provided on the impact of the desired polarization state, beam direction, and array size. The discussion points out that the array size is a crucial factor for the performance of the proposed design.
The remainder of this paper is organized as follows. Section II introduces the general design of the PCRPA and its beamformer architecture. Section III proposes the methods of polarization pattern synthesis based on the PRCA and verifies them through simulations. Section IV analyzes the parameter sensitivity of pattern synthesis, including the parameters of polarization state, array size, and beam position. Section V presents the experiment results to verify further the effectiveness of the proposed architecture and synthesis methods concerning PPA radars, then discusses the applicable scenarios. Section VI concludes the entire paper.
Notations: Some important symbols and notational definitions used in this article are given in Table I. In this paper, boldface upper-case letters denote matrices, boldface lower-case letters denote column vectors, and standard lowercase letters denote scalars.
TABLE I: Notation
Symbol Description
$\theta$ Elevation angle
$\phi$ Azimuth angle
$\gamma$ Auxiliary polarization angle
$\eta$ Polarization phase angle
H Horizontal
V Vertical
CoP Co-polarization
XP Cross-polarization
PPA Polarimetric phased array
CPPA Conventional polarimetric phased array
PCRPA Polarization coding reconfigurable phased array
PSL Peak sidelobe level
XPL Cross-polarization level
AEP Active element pattern
II General Architecture of The PCRPA
Figure 1: The illustration of the conventional and the proposed polarimetric phased array architectures. (a) Architecture of the conventional polarimetric phased array, where two polarization ports of each element are connected to independent channels. (b) Architecture of the proposed polarimetric phased array with polarization reconfigurable elements. (c) Schematic diagram of the proposed array, which can produce dual-polarized or arbitrarily polarized beams by controlling the polarization states and waveforms of elements.
The architecture designs of the conventional PPA and the proposed PCRPA are illustrated in Fig. 1a and 1b. Although both diagrams adopt uniform $N$-element linear arrays with the spacing $d$ for simplicity, the proposed architecture can also be utilized in various antennas, including planar and conformal arrays. The elements for the conventional PPA are dual-polarized antennas consisting of two orthogonal polarization ports (usually H and V polarization), represented by the red and blue rectangles in Fig. 1a. Regardless of the subarray design, each port is connected to an independent T/R channel, corresponding to the independent weighting $a_{ih}\phi_{ih}$ or $a_{iv}\phi_{iv}$ and waveform H or V. In this case, the electric field of beams formed by H and V ports can be expressed as:
$\left\{{\begin{array}[]{l}{\boldsymbol{E}_{h}\left({\theta,\phi}\right)=\sum_{i=1}^{N}\boldsymbol{e}_{ih}\left({\theta,\phi}\right)}\\ {\boldsymbol{E}_{v}\left({\theta,\phi}\right)=\sum_{i=1}^{N}\boldsymbol{e}_{iv}\left({\theta,\phi}\right)}\end{array}}\right$. (1)
where $\boldsymbol{e}_{ih}(\theta,\phi)$ and $\boldsymbol{e}_{iv}(\theta,\phi)$ denote the electric field vectors of the $i$-th H and V ports, and the effects of amplitude weights $a_{ih}$, $a_{iv}$ and phase weights $\phi_{ih}$, $\phi_{iv}$ have already been included.
For the proposed PCRPA architecture, the elements are polarization reconfigurable and represented by the spliced rectangles in Fig. 1b. The colored and transparent halves of the rectangles indicate polarization excited and unexcited, respectively. These elements are each connected to a single T/R channel, corresponding to the independent weighting $a_{i}\phi_{i}$ and waveform H or V. Compared to the conventional PPA, the number of T/R channels here is reduced to half. The PCRPA can form multiple polarized beams by controlling the coding states of elements, where the beam can be expressed as:
$\boldsymbol{E}_{j}(\theta,\phi)=\sum_{i=1}^{N_{jh}}\boldsymbol{e}_{I^{jh}_{i}h}(\theta,\phi)+\sum_{i=1}^{N_{jv}}\boldsymbol{e}_{I^{jv}_{i}v}(\theta,\phi)$ (2)
where $\boldsymbol{E}_{j}(\theta,\phi)$ is the electric field of the $j$-th beam, $\boldsymbol{I}^{jh}$ and $\boldsymbol{I}^{jv}$ are vectors containing the element indices coded by H and V polarization for the $j$-th beam, $N_{jh}$ and $N_{jv}$ are the number of these elements. Here, vectors $\boldsymbol{I}^{jh}$ and $\boldsymbol{I}^{jv}$ should satisfy the constraints followed:
$\left\{{\begin{array}[]{l}{1\leq I^{jA}_{i}\leq N,\>I^{jA}_{i}\in\mathbb{Z},i=1,2,\cdots,N_{jA}}\\ {I^{jA}_{m}\neq I^{jA}_{n},\>\forall m,n\in\{1,2,\cdots,N_{jA}\},m\neq n}\\ {\boldsymbol{I^{jh}\cap\boldsymbol{I}^{jv}=\emptyset}}\end{array}}\right$. (3)
where $A\in\{h,v\}$ represents the polarization port.
Therefore, it can be summarized from Fig. 1b and Equation (2)-(3) that beam patterns of the PCRPA are synthesized by 3 groups of variable: the amplitude and phase weights, the waveform code and the polarization code. The latter two groups exist only in the proposed architecture rather than the conventional PPA. By adjusting these variables in real-time, the PCRPA can form multiple electronically scanned beams with any desired polarization. Considering the practical scenarios in radar applications, this paper focuses on two problems: arbitrarily polarized and dual-polarized pattern synthesis, shown in Fig. 1c. These beam patterns of high performance can meet the requirements of polarization measurements and anti-interference.
Since the proposed PCRPA architecture includes a two-level coding process, it can be realized by the beamformer designed with two-level RF switches shown in Fig. 2. Here, RF switches are utilized to connect the array antenna, T/R modules, and transceivers sequentially. To highlight the architectural features, the radar transmitters and receivers are schematically represented in the diagram. Regarding digitization levels: The current design reflects the active phased array architecture; Subarray digital architecture can be achieved by incorporating multiple digital transceivers after the T/R modules; For fully digital architectures, the second-level switches become redundant and can be eliminated[37]. This is because the waveforms can be conveniently altered by direct digital synthesizers (DDS) in digital transceivers at the element level. Besides, the combination of dual-polarized antennas and RF switches can also be replaced by polarization reconfigurable antennas with single ports. However, the latter may not tolerate high transmitting power due to components like PIN diodes.
Figure 2: The beamformer design of the proposed phased array architecture for active arrays or digitized subarrays, where RF switches are utilized between T/R modules, array antennas, and transceivers.
III Pattern Synthesis for The PCRPA
III-A Problem Formulation
Consider a conventional polarimetric phased array antenna with $N$ elements, the vector of its active element patterns (AEPs) can be expressed as:
$\boldsymbol{g}_{AB}(\theta,\phi)=\left[g_{1,AB}(\theta,\phi),...,g_{N,AB}(\theta,\phi)\right]^{T}$ (4)
where $A,B\in\{h,v\}$ represent the polarization of the electric field component and the antenna port, respectively.
The beam patterns are formed for the active phased arrays by adding each AEP with independent weights. In this case, the array pattern can be expressed as[38]:
$\left\{{\begin{array}[]{l}{{f_{h}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{g}_{hh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{g}_{hv}(\theta,\phi)}\\ {{f_{v}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{g}_{vh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{g}_{vv}(\theta,\phi)}\end{array}}\right$. (5)
where $\boldsymbol{w}_{h}$ and $\boldsymbol{w}_{v}$ denote the weight vectors for H and V ports of the array. These vectors can be generally adjusted freely, depending on the amplitudes and phases, to form desired beam patterns. Due to space limitations, this paper only addresses the transmitting problem, which allows phase-only synthesis model below:
$\boldsymbol{w}_{h}=\boldsymbol{w}_{v}=\text{exp}(jk\boldsymbol{P}^{T}\boldsymbol{r}_{0})$ (6)
where $k=2\pi/\lambda$, $\boldsymbol{r}_{0}=\left[\sin{\theta_{0}}\cos{\phi_{0}},\sin{\theta_{0}}\sin{\phi_{0}},\cos{\theta_{0}}\right]^{T}$, and $\boldsymbol{P}$ represents the position of elements:
[…]
According to the architecture shown in Fig. 1b, the synthesis of PCRPA patterns depends on variables of the weight and code vectors for antenna elements. The coding process can be expressed by the polarization coding vectors $\boldsymbol{x}_{h}$ and $\boldsymbol{x}_{v}$. The vectors should satisfy:
$\left\{{\begin{array}[]{l}{x_{h},x_{v}\in\{0,1\}}\\ {\boldsymbol{x}_{h}+\boldsymbol{x}_{v}\leq\boldsymbol{1}}\end{array}}\right$. (8)
For clear clarification, define code matrix as below:
$\left\{{\begin{array}[]{l}{\boldsymbol{X}_{h}=\text{diag}(\boldsymbol{x}_{h})}\\ {\boldsymbol{X}_{v}=\text{diag}(\boldsymbol{x}_{v})}\end{array}}\right$. (9)
For any beam synthesized by the PCRPA, its pattern can be then expressed as:
$\left\{{\begin{array}[]{l}{{f_{h}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{X}_{h}\boldsymbol{g}_{hh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{X}_{v}\boldsymbol{g}_{hv}(\theta,\phi)}\\ {{f_{v}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{X}_{h}\boldsymbol{g}_{vh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{X}_{v}\boldsymbol{g}_{vv}(\theta,\phi)}\end{array}}\right$. (10)
III-B Arbitrarily Polarized Pattern Synthesis
The subsequent part introduces the arbitrarily polarized pattern synthesis method procedure, with a detailed flowchart illustrated in Fig. 3.
Figure 3: Flowchart of the arbitrarily polarized pattern synthesis method.
For arbitrarily polarized pattern synthesis problems, define the desired polarization basis according to Ludwig’s polarization definition II[39]:
$\left\{\begin{array}[]{l}{\boldsymbol{e}_{co}=\left[\cos{\gamma,\sin{\gamma}e^{j\eta}}\right]^{T}}\\ {\boldsymbol{e}_{cr}=\left[-\sin{\gamma,\cos{\gamma}e^{-j\eta}}\right]^{T}}\\ {\boldsymbol{e}_{co}^{H}\boldsymbol{e}_{cr}=0}\end{array}\right$. (11)
where $\gamma\in\left[0^{\circ},90^{\circ}\right]$ and $\eta\in\left[-180^{\circ},180^{\circ}\right)$ represent the parameters of the auxiliary polarization angle and the polarization phase difference. The vectors of $\boldsymbol{e}_{co}$ and $\boldsymbol{{e}_{cr}}$ denote the unit electric field vectors of the co-polarization and cross-polarization components.
According to the conservation of electric fields, the following equation holds:
$\left[\boldsymbol{e}_{co},\boldsymbol{e}_{cr}\right]\left[\begin{array}[]{l}{f_{co}(\theta,\phi)}\\ {f_{cr}(\theta,\phi)}\end{array}\right]=\left[\boldsymbol{e}_{h},\boldsymbol{e}_{v}\right]\left[\begin{array}[]{l}{f_{h}(\theta,\phi)}\\ {f_{v}(\theta,\phi)}\end{array}\right]$ (12)
where $\boldsymbol{e}_{h}=\left[1,0\right]$ and $\boldsymbol{e}_{v}=\left[0,1\right]$ represent the unit vectors of the H and V polarization fields.
Let the co-polarization component $f_{co}(\theta,\phi)=1$ and cross-polarization component $f_{cr}(\theta,\phi)=0$, according to Equation (12), the vector decomposition results of the desired polarization state under the standard H and V basis can be defined as:
$\left[\begin{array}[]{l}{u_{h}}\\ {u_{v}}\end{array}\right]$ $\triangleq\left[\begin{array}[]{l}{f_{h}(\theta,\phi)}\\ {f_{v}(\theta,\phi)}\end{array}\right]$ (13)
$=\left[\boldsymbol{e}_{h},\boldsymbol{e}_{v}\right]^{-1}\boldsymbol{e}_{co}$
$=\left[\begin{array}[]{l}{\cos(\gamma)}\\ {\sin(\gamma)e^{j\eta}}\end{array}\right]$
Based on Equation (13), the conventional PPAs can synthesize patterns of arbitrarily polarized beams by directly multiplying the corresponding weights $\cos{\gamma}$ and $\sin{\gamma}e^{-j\eta}$ into the polarization channels H and V[36]. However, this method ignores the cross-polarization components radiated at the H and V ports. If they are considered, there will be:
$f_{cr}(\theta,\phi)=$ $\sin(\gamma)\cos(\gamma)\left[-f_{hh}(\theta,\phi)+f_{vv}(\theta,\phi)\right]$ (14)
$+\cos^{2}(\gamma)e^{-j\eta}f_{vh}(\theta,\phi)-\sin^{2}(\gamma)e^{j\eta}f_{hv}(\theta,\phi)$
Even if $f_{hh}(\theta,\phi)$ and $f_{vv}(\theta,\phi)$ are highly matched, due to the influence of the latter two terms, the cross-polarization component of the beam $f_{co}(\theta,\phi)$ is difficult to suppress. Especially when the scanning angle increases, the polarization purity of the beam will decrease rapidly.
In this case, this paper proposes a decomposition model based on antenna AEPs. The idea is shown in the second part of Fig. 3. The green arrow represents the desired polarization state of the beam pattern. It can be decomposed according to the polarization basis represented by the red and blue arrows. The vectors of the basis are defined by the AEPs of an embedded element taken from the array center:
$\left\{{\begin{array}[]{l}{\boldsymbol{\tilde{e}}_{h}=\boldsymbol{g}_{h}(\theta_{0},\phi_{0})/\left\|\boldsymbol{g}_{h}(\theta_{0},\phi_{0})\right\|}\\ {\boldsymbol{\tilde{e}}_{v}=\boldsymbol{g}_{v}(\theta_{0},\phi_{0})/\left\|\boldsymbol{g}_{v}(\theta_{0},\phi_{0})\right\|}\end{array}}\right$. (15)
where
[…]
And the decomposition process can be expressed as:
$\left[\begin{array}[]{l}{\tilde{u}_{h}}\\ {\tilde{u}_{v}}\end{array}\right]=\left[\boldsymbol{\tilde{e}}_{h},\boldsymbol{\tilde{e}}_{v}\right]^{-1}\boldsymbol{e}_{co}$ (17)
where $\tilde{u}_{h}$ and $\tilde{u}_{v}$ are complex numbers that can measure the contribution of each port to the desired polarization.
Since this paper focuses on the phase-only synthesis, the amplitude of weights $\tilde{u}_{h}$ and $\tilde{u}_{v}$ can be used to synthesize the number of elements excited by H and V ports:
$\left\{{\begin{array}[]{l}{n_{h}\triangleq\lVert\boldsymbol{x}_{h}\rVert_{1}=\lceil{\frac{{\left|\tilde{u}_{h}\right|}}{{\left|\tilde{u}_{h}\right|+\left|\tilde{u}_{v}\right|}}N}\rceil}\\ {n_{v}\triangleq\lVert\boldsymbol{x}_{v}\rVert_{1}=N-n_{h}}\end{array}}\right$. (18)
In addition, a compensation phase is applied to the channels of V ports to ensure the desired polarization, which is defined:
$\beta=\angle\left({\frac{\tilde{u}_{v}}{\tilde{u}_{h}}}\right)$ (19)
This phase component is directly multiplied by the weight vector of $\boldsymbol{w}_{v}$ given by Equation (6):
$\boldsymbol{w}_{v}=\text{exp}\left[j(k\boldsymbol{P}^{T}\boldsymbol{r}_{0}+\beta)\right]$ (20)
The array configuration of $\boldsymbol{x}_{h}$ and $\boldsymbol{x}_{v}$ is yet to be synthesized, which strongly relates to the sidelobe levels of the patterns. To fully investigate the pattern characteristics under the PCRPA model, Fig. 4 gives a histogram of the Monte Carlo simulation results concerning this method. The array size is $16\times 16$, with the desired polarization of V polarization steered at $(120^{\circ},15^{\circ})$. The red dot line marks the PSL of the CPPA under the same circumstances using the weight vector $\boldsymbol{w}_{v}$ given in Equation (6). The distribution fitting represented by the red solid line shows that 29.74% of the samples have PSLs lower than the conventional architecture, and the worst case is only 1.63 dB higher. This indicates a relatively concentrated distribution, which is convenient to optimize.
Figure 4: Histogram of the Monte Carlo simulation results concerning arbitrarily polarized pattern synthesis.
The whole method can be summarized step by step, given in Algorithm 1. Although Equation (18) can uniquely calculate the number of H and V elements, this method provides an adjustment mechanism realized by Steps 4 and 7. This is because the AEPs of embedded elements may not fully represent the array polarization characteristics due to mutual coupling and other effects. For each $n_{h}$ and $n_{v}$, a maximum generation is set. If the iteration reaches the maximum generation but the patterns do not meet the requirements, the method will adjust $n_{h}$ and $n_{v}$. The adjustment rules can be flexible, for example, reducing one H element each time. The random number generator in Step 5 is chosen according to simulation results given by Fig. 4, which can also be replaced by other optimization algorithms if needed.
Algorithm 1 The proposed arbitrarily polarized pattern synthesis method for the PCRPA.
1: Input the data of active element patterns $\boldsymbol{g}_{AB}(\theta,\phi)$, obtained from either simulations or experiments. Initialize the beam direction $(\theta_{0},\phi_{0})$, desired polarization state $(\gamma_{0},\eta_{0})$, the pattern requirements $\Gamma_{PSL}$ and $\Gamma_{XPL}$, and the mainlobe region $\Omega_{mb}$. Set the termination condition of the run time or the maximum number of generations.
2: Calculate the decomposition results of the desired polarization state based on the AEP model given by Equation (15)-(19).
3: Calculate the number of H and V polarized elements $n_{h}$, $n_{v}$, and the compensation phase $\beta$ according to the polarization basis decomposition results.
4: Check whether the number of H and V polarized elements is reasonable. If yes, proceed to Step 5; otherwise, stop the synthesis procedure and return the failure prompt.
5: Generate the array configuration $\boldsymbol{x}_{h}$, $\boldsymbol{x}_{v}$ according to $n_{h}$, $n_{v}$ using random number generators. Calculate the patterns $f_{co}(\theta,\phi)$, $f_{cr}(\theta,\phi)$ using Equation (10) and (12).
6: Check if the PSL and XPL of the patterns $f_{co}(\theta,\phi)$, $f_{cr}(\theta,\phi)$ meet the requirements of $\Gamma_{PSL}$, $\Gamma_{XPL}$. If yes, exit the procedure and return the synthesis results; otherwise, proceed to Step 7.
7: Check if the iteration has reached the maximum generation. If yes, adjust $n_{h}$, $n_{v}$ and repeat Steps 4 to 6; otherwise, repeat Steps 5 to 6.
III-C Dual-polarized Pattern Synthesis
For the measurement applications of the STSR mode, a pair of orthogonally polarized beams is required simultaneously. Therefore, all the elements are split into two groups to synthesize the beams. This can be represented by the waveform coding vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, where $x_{1},x_{2}\in\{0,1\}$ and $x_{i,j}=1$ means the $i$-th element belongs to the $j$-th beam. Take the beam $1$ as an example, Equation (10) can be transferred to:
$\left\{{\begin{array}[]{l}{{f_{1h}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{X}_{1h}\boldsymbol{g}_{hh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{X}_{1v}\boldsymbol{g}_{hv}(\theta,\phi)}\\ {{f_{1v}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{X}_{1h}\boldsymbol{g}_{vh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{X}_{1v}\boldsymbol{g}_{vv}(\theta,\phi)}\end{array}}\right$. (21)
where
$\left\{{\begin{array}[]{l}{\boldsymbol{X}_{1h}=\text{diag}(\boldsymbol{x}_{1h})}\\ {\boldsymbol{X}_{1v}=\text{diag}(\boldsymbol{x}_{1v})}\\ {\boldsymbol{x}_{1h}=\boldsymbol{x}_{1}\odot\boldsymbol{x}_{h}}\\ {\boldsymbol{x}_{1v}=\boldsymbol{x}_{1}\odot\boldsymbol{x}_{v}}\end{array}}\right$. (22)
In this case, the synthesis of dual-polarized beam patterns is decided by variables of the beam coding vectors $\boldsymbol{x}_{1h}$, $\boldsymbol{x}_{1v}$, $\boldsymbol{x}_{2h}$, and $\boldsymbol{x}_{2v}$, which are defined by the Hadamard product of the polarization and waveform coding vectors.
According to [40], the calibration methods can conveniently mitigate the cross-polarization components of these dual-polarized beams. Therefore, it is reasonable to assume that no cross-polarized elements are involved in each beam, which can be expressed as:
$\left\{{\begin{array}[]{l}{\boldsymbol{x}_{1}=\boldsymbol{x}_{h}}\\ {\boldsymbol{x}_{2}=\boldsymbol{x}_{v}}\end{array}}\right$. (23)
Since Equation (23) holds, the variables $\boldsymbol{x}_{1v}$ and $\boldsymbol{x}_{2h}$ are equal to zero, from which can infer that:
$\boldsymbol{x}_{1h}+\boldsymbol{x}_{2v}=\boldsymbol{1}$ (24)
where $\text{sum}(\boldsymbol{x}_{1h})=\text{sum}(\boldsymbol{x}_{2v})=N/2$ to ensure good match of the co-polar beams.
A Monte Carlo simulation given by Fig. 5 has been carried out according to Equation (24) to investigate the PSL distribution under the dual-polarized pattern synthesis model. The beampoint is set at $(120^{\circ},15^{\circ})$, with the PSL taking the maximum of the H and V beams. The red dot line marks the value calculated under the CPPA using Equation (6). Unlike the phenomenon revealed by Fig. 4, this histogram shows a more dispersed distribution, with the worst PSL 5 dB higher than the CPPA. In addition, the fitting curve marked by the red solid line indicates that only 0.92% of the configurations have PSLs that are no worse than that of the CPPA.
Figure 5: Histogram of the Monte Carlo simulation results concerning dual-polarized pattern synthesis.
Therefore, an optimized model with a central symmetry constraint is introduced to derive low PSL results. This model can be solved using binary genetic algorithm (BGA) methods and has been discussed in detail in previous work by the authors[35]. The central symmetry constraint means that two configurations forming the H and V beams are centrally symmetric shown in Fig. 6. In this case, the array factors produced by these configurations will have the same power patterns, which have been theoretically demonstrated in [35].
Figure 6: Schematic diagram of the central symmetry constraint.
This property ensures the optimization model derives highly matched co-polar beam patterns. The constraint can be expressed by the following equation:
$\boldsymbol{X}_{2v}=\boldsymbol{1}-\text{rot180}(\boldsymbol{X}_{1h})$ (25)
where $\text{rot180(}\cdot\text{)}$ means to rotate a matrix 180 degrees.
From the perspective of matrix operations, Equation (25) can be further represented in vector form using the code vectors $\boldsymbol{x}_{1h}$ and $\boldsymbol{x}_{2v}$:
[…]
where $\boldsymbol{T}$ is defined as the rotational transformation matrix. Each row or column of $\boldsymbol{T}$ contains only one element equal to one and the others equal to zero, and $T_{ij}=1$ means that the $i$-th and the $j$-th elements are centrally symmetric in the grid. In this paper, the 1-elements are arranged on the subdiagonal of the matrix.
The pattern synthesis problem can be then formulated as a minimax problem for minimizing the PSL of $f_{1h}\left(\theta,\phi\right)$ and $f_{2v}\left(\theta,\phi\right)$ by determining the coding of elements in the binary vector $\boldsymbol{x}_{u}$, which is given by
[…]
This nonlinear integer programming problem tends to be solved by stochastic optimization algorithms such as BGA, binary PSO (BPSO), and binary differential evolution (BDE). In this paper, the BGA algorithm is adopted. The whole procedure of the synthesis method is shown in Algorithm 2 and explained step by step.
Algorithm 2 The proposed dual-polarized pattern synthesis method for the PCRPA.
1: Input the data of active element patterns $\boldsymbol{g}_{AB}(\theta,\phi)$, obtained from either simulations or experiments. Initialize the beam direction $(\theta_{0},\phi_{0})$ and the sidelobe region $\Omega_{sl}$. Set the termination condition of the run time or the maximum number of generations.
2: Optimize the variable $\boldsymbol{x}_{u}$ using the BGA algorithm according to the model given by Equation (27). Stop the process if it reaches the termination condition.
3: Calculate the optimized configuration results according to Equation (26).
4: Calibrate the dual-polarized patterns according to the method mentioned in [40].
III-D Simulation Results
Input data from the HFSS model of a $16\times 16$-unit dual-polarized microstrip array antenna is adopted to demonstrate the performance of the proposed architecture and methods. The simulations are carried out for vertically, arbitrarily, and dual-polarized patterns concerning various radar applications.
III-D1 Vertically Polarized Pattern
This simulation considers a V polarization pattern with the beam direction at $(120^{\circ},15^{\circ})$. These beams with pure V or H polarization are standard in polarization-measuring applications of ATSR modes. According to the initial conditions and the method given in Algorithm 1, the results of the element polarization configuration are shown in Fig. 7. In this configuration, 17 out of 256 elements, which account for 6.64%, are excited by the H-pol ports to eliminate the cross-polarization components of the synthesized beam. Their positions are generated based on the proposed method.
Fig. 8 presents the 3-D normalized power patterns of different array architectures or methods under the same data input and initial conditions. The calibration method aimed at the ATSR mode is adopted from [40]. Compared to the CPPA, PCRPA exhibits a noticeable improvement in XP component mitigation, with the value of the cross-polarization level XPL decreasing from -22.93 dB to -59.68 dB, as shown in Fig. 8b and 8f. This is comparable to the effects of calibration, as shown in Fig. 8d, where the XPL is -65.31 dB. Meanwhile, the CoP pattern of PCRPA indicates only an even better level in the PSL compared to conventional PPA, as shown in Fig. 8a, 8c, and 8e.
Figure 7: H-pol and V-pol element distribution of a $16\times 16$-unit planar array for the required polarization state $(90^{\circ},0^{\circ})$ and beam direction $(120^{\circ},15^{\circ})$. The blue and red squares represent H-pol and V-pol elements, respectively.
Figure 8: The 3-D normalized power patterns of the simulated $16\times 16$-unit microstrip antenna array, where V polarization is the co-polarization and the beam direction is $(120^{\circ},15^{\circ})$. The red dashed line helps indicate the pointing direction of the array beam. (a) The CoP pattern before calibration of the CPPA. (b) The XP pattern before calibration of the CPPA. (c) The CoP pattern after calibration of the CPPA. (d) The XP pattern after calibration of the CPPA. (e) The CoP pattern of the proposed PCRPA. (f) The XP pattern of the proposed PCRPA.
TABLE II: Comparison of the key parameters of the vertically polarized pattern synthesis
Group CPPA, uncalibrated CPPA, calibrated PCRPA
XPL/dB -22.93 -65.31 -59.68
PSL/dB -12.75 -12.77 -13.39
Pt/dB 0 -0.01 -0.55
Dir/dBi 28.37 28.37 28.15
To clarify in detail, Table III presents the specific parameter values of the three patterns. Here, $P_{t}$ represents the transmission power at the beam direction normalized with the value of the uncalibrated beam under CPPA; $D$ denotes the directivity of the array pattern. As indicated by the data, since some elements are synthesized to suppress XP components, the integrating sidelobes of both the CoP and XP patterns of PCRPA have risen, leading to a loss of power and directivity. The results show that $P_{t}$ and $D$ of PCRPA are reduced by 0.55 dB and 0.22 dB, respectively.
III-D2 Arbitrarily Polarized Pattern
This simulation considers an arbitrarily polarized pattern with the same beam direction as in the above section. Such requirements tend to appear in many anti-interference applications, where the beam polarization is orthogonal to the interference signal polarization to suppress its received power. Here, the desired polarization state is set to $(30^{\circ},60^{\circ})$. According to Equation (13), the 2-D normalized power patterns of the CPPA are given by Fig. 9. The XPL of this beam is at a relatively high level of -21.25 dB, which may affect the anti-interference processing.
Figure 9: The azimuth cut plane of the 2-D normalized power pattern of the arbitrarily polarized pattern synthesis under the CPPA architecture, with the array size of $16\times 16$, co-polarization state at $(30^{\circ},60^{\circ})$ and the beam direction at $(120^{\circ},15^{\circ})$.
Fig. 10 presents three groups of PCRPA results obtained by different pattern synthesis methods, including the proposed, random, and BGA methods. Table III lists the key parameters of these results and the CPPA in detail. Here, $\text{XPL}_{\text{m}}$ denotes the normalized peak level of the XP pattern across the scan principal planes. The difference between these three methods lies in Step 4 to Step 6 of the algorithm given in subsection III-B. The random method obtains the array configuration by performing the random number generator only once, where the XPL and PSL constraints are ineffective. The BGA method obtains the array configuration by optimizing the coding vectors $\boldsymbol{x}_{h}$ and $\boldsymbol{x}_{v}$ using a binary genetic algorithm. Computations were performed on a PC with an Intel Core i7-10510U central processing unit (CPU) at 2.30 GHz and 16 GB RAM.
Figure 10: Comparison of the arbitrarily polarized pattern synthesis results for the proposed architecture using different methods. (a) Elevation plane using the proposed method. (b) Elevation plane using the random method. (c) Elevation plane using the GA method. (d) Azimuth plane using the proposed method. (e) Azimuth plane using the random method. (f) Azimuth plane using the GA method. (g) Array configuration using the proposed method. (h) Array configuration using the random method. (i) Array configuration using the GA method.
TABLE III: Comparison of key parameters of the arbitrarily polarized pattern synthesis
Group CPPA PCRPA
proposed random BGA
XPL/dB -24.43 -49.13 -49.13 -49.05
$\text{XPL}_{\text{m}}$/dB -24.15 -20.29 -8.51 -20.61
PSL/dB -12.75 -13.38 -12.63 -13.93
Run time/s \ 5.19 \ 1206.95
Considering the comparison of XPL and PSL between CPPA and PCRPA, the simulation results here are consistent with the V polarization simulations. The advantage of PCRPA’s cross-polarization suppression still exists, and so does the decrease of the peak sidelobe, which is at a value of -13.38 dB. Besides, among the 3 groups of PCRPA results, the XP patterns of the random method show an excessively high level at -8.71 dB within the main lobe area, given in Fig. 10b and 10e. This excessively high sidelobe of XP patterns is eliminated in the other 2 groups due to the optimization of element configurations, as shown in Fig. 10a, 10c, 10d, and 10f. Furthermore, the proposed method and the BGA method’s pattern synthesis results indicate comparable performance, with the proposed method’s PSL being 0.55 dB higher. However, its run time is also significantly shorter, only 5.19 seconds compared with 1206.95 seconds using the BGA method according to Table III, which is an obvious efficiency advantage. The simulation results presented above show that the proposed method can achieve a compromise between performance and efficiency in the PCRPA architecture.
III-D3 Dual-polarized Patterns
Figure 11: H-pol and V-pol element distribution of a $16\times 16$-unit planar array for dual-polarized beams with beam direction $(120^{\circ},15^{\circ})$. The blue squares and red diamonds represent H-pol and V-pol elements, respectively.
Figure 12: Comparison of the dual-polarized pattern synthesis results for the conventional and proposed architecture using different methods. (a) Elevation plane of the CPPA. (b) Elevation plane of the PCRPA using the proposed method. (c) Elevation plane of the PCRPA using the random method. (d) Azimuth plane of the CPPA. (e) Azimuth plane of the PCRPA using the proposed method. (f) Azimuth plane of the PCRPA using the random method.
TABLE IV: Comparison of different groups
Group CPPA PCRPA
proposed random
XPL/dB H -56.89 -56.76 -56.86
V -57.80 -57.87 -57.84
PSL/dB H -12.74 -12.73 -11.70
V -12.77 -12.76 -8.95
$P_{t}$/dB H 0 -6.02 -6.02
V -0.02 -6.04 -6.04
DD/dBi H 28.37 26.12 25.86
V 28.37 26.12 26.07
ME/dB -19.57 -19.59 10.36
This simulation considers a pair of H and V polarization beams with the same beam direction as above, which are commonly required in polarization measuring applications of the STSR mode. According to the proposed synthesis method given by Equation (27) and Algorithm 2, the entire array is split into two groups, each fed with the corresponding waveform. To quantitatively evaluate the matching performance of the main lobes in dual-polarized beams, a new parameter called matching error (ME) is introduced, defined as[41]:
[…]
where $f_{1h,k}$ and $f_{2v_{k}}$ represent the $k$-th sampling points within the main lobe of patterns $f_{1h}(\theta,\phi)$ and $f_{2v}(\theta,\phi)$.
Fig. 12 compares three groups of dual polarization patterns obtained under different configurations: the CPPA, the optimized PCRPA, and a bad case of PCRPA. All the patterns have been processed by polarization calibration methods. Fig. 11 shows the configuration of the optimized PCRPA. Markers of squares and diamonds represent the H and V polarization excited elements to distinguish the coding of waveforms. The bad case configuration is obtained without the central symmetry constraint proposed in this paper, simply using a random number generator instead. The optimized PCRPA performs well on the matched co-polar beams shown in Fig. 12b and 12e. Its ME value is very close to that of conventional PPA and 29.96 dB better than that of the bad case, which verifies the effectiveness of the constraint. The same phenomenon also applies to the PSL parameter detailed in Table IV. In addition to the above advantages, Table IV shows that the PCRPA has a significant loss problem on power and directivity at around 6 dB and 2.2 dB, respectively, compared to the CPPA. This may limit the maximum radar range under STSR mode.
IV Parameter Sensitivity Analysis
Figure 13: Pattern performance of the proposed and conventional polarimetric phased arrays under different desired polarization states. (a) PSL and XPL under different $\gamma$. (b) Pt and Dir under different $\gamma$. (c) PSL and XPL under different $\eta$. (d) Pt and Dir under different $\eta$.
The simulation results have verified the effectiveness of the proposed PCRPA architecture and synthesis methods. This section explores the sensitivity of pattern performance to various input parameters, including the polarization state, beam direction, and array size.
IV-A Polarization State
The desired polarization state affects the pattern performance by modifying the polarization vector’s decomposition results given by Equation (17), and further changing element numbers $n_{h}$, $n_{v}$, and compensation phase $\beta$.
Fig. 13 presents the proposed PCRPA and the CPPA pattern performance concerning different desired polarization states. The results are obtained concerning the synthesis of arbitrarily polarized patterns. The polarization state parameters are adjusted by varying either $\gamma$ (with $\eta$ fixed at $60^{\circ}$) or $\eta$ (with $\gamma$ fixed at $30^{\circ}$). First, compared to the curves shown in Fig. 13c and 13d, the PCRPA curves in Fig. 13a and 13b exhibit more pronounced fluctuations, indicating a stronger sensitivity to $\gamma$ than $\eta$. Then, considering the impact of $\gamma$, the XPL is consistently suppressed below -40 dB and exhibits little correlation with the polarization state. In contrast, the PSL, directivity, and power curves show a more regular trend that changes with $\gamma$. When $\gamma$ approaches $45^{\circ}$, the PSL and directivity become close to the CPPA values, but the power loss increases to a maximum of around 3 dB. This is because the number of H- and V-polarized elements $n_{h}$ and $n_{v}$ tend to be equal, resulting in similar array factors to those of the CPPA. Meanwhile, a significant portion of power is canceled during the vector composing process.
These can lead to the following conclusions. For a given array size and beam direction:
  • The XPL in PCRPA pattern synthesis is almost insensitive to the desired polarization state, while the PSL, directivity, and radiated power are sensitive to the auxiliary polarization angle $\gamma$.
  • The PCRPA exhibits lower PSL and directivity compared to the CPPA. Their performance difference gradually diminishes as $\gamma$ approaches $45^{\circ}$, until the characteristics become nearly comparable.
  • The PCRPA exhibits a certain level of power loss compared to the CPPA. The loss magnitude progressively increases as $\gamma$ approaches $45^{\circ}$, with the maximum loss reaching approximately 3 dB.
IV-B Beam Direction
Figure 14: Pattern performance of the proposed and conventional polarimetric phased arrays under different beam directions. In the legend, ’1’ represents the single-polarized beam, and ’2’ denotes the dual-polarized beam. (a) PSL curves. (b) XPL curves. (c) Directivity curves.
Figure 15: Pattern performance of the proposed and conventional polarimetric phased arrays under different array sizes. In the legend, ’1’ represents the single-polarized beam, and ’2’ denotes the dual-polarized beam. (a) PSL curves. (b) XPL curves. (c) Directivity curves.
The beam direction affects the pattern performance by determining the specific basis vector sets selected through Equation (15), as well as the element pattern data and weight vectors used in Equation (21).
Fig. 14 shows the performance of both single- and dual-polarized pattern synthesis of the PCRPA and the CPPA concerning different beam directions. The azimuth angle $\phi$ is fixed at $15^{\circ}$ while the elevation angle $\theta$ ranges from $30^{\circ}$ to $150^{\circ}$ with a step of $15^{\circ}$. The polarization state of the single-polarized pattern is fixed at $(30^{\circ},60^{\circ})$. Figs. 14a and 14c reveal the same degradation trends of PSL and directivity for both architectures concerning single- and dual-polarized cases with increasing scanning angles. Additionally, the performance gaps between the PCRPA and the CPPA remain stable throughout the scanning range. In Fig. 14b, the XPL of the single-polarized pattern appears insensitive to scanning angle variations. In contrast, the dual-beam XPL demonstrates an increasing trend as the scanning angle $\Delta\theta$ enlarges. When $\Delta\theta\leq 30^{\circ}$, the dual-beam XPL is lower than the single-beam by 10-20 dB; however, they converge to comparable values when $\Delta\theta>30^{\circ}$. This owes to the different XPL suppression mechanisms: the dual-beam employs the calibration matrix with continuously adjustable coefficients, while the single-beam relies on controlling the ratio of H- and V-polarized elements, where the element numbers are discrete. In addition, beam pointing error at large scanning angles induces a mismatch in the calibration matrix, deteriorating the XPL in dual-beam synthesis.
These can lead to the following conclusions. For a given array size and polarization state:
  • All pattern performance of the PCRPA, except the XPL of the single-polarized beam, exhibits sensitivity to beam scanning, with progressive degradation observed as the scanning angle increases.
  • The gaps in PSL and directivity between the PCRPA and CPPA remain almost stable throughout the scanning range.
  • The XPL of the PCRPA is always suppressed at no more than -40 dB, where the dual beam pattern synthesis can exhibit excessively low levels at -60 dB near the broadside.
IV-C Array Size
The array size affects the pattern performance by determining the specific number of H- and V-polarized elements in Equation (18), and the dimensionality of the optimization model given by Equation (27).
As shown in Fig. 15, all radiation pattern parameters exhibit evident improvements with the array size, except the XPL of dual-beam for the PCRPA. Notably, the XPL of dual-beam PCRPA consistently maintains -60 dB-level stability and demonstrates insensitivity to the array size. This is because its XP suppression performance is determined by the estimation accuracy of the polarization calibration matrix rather than the array size. Particularly, both the PSL of dual-beam synthesis and the XPL of single-beam synthesis show remarkable array size dependence in small-to-medium arrays ($\leq 24\times 24$): The performance enhancement in dual-beam synthesis stems from effective dimensionality expansion of the optimization model. In contrast, for the single-beam case, the improvement arises from the quantization characteristics of the polarization angle $\gamma$, given by the equation below:
$\hat{\gamma}=\arctan(\frac{n_{v}}{n_{h}})$ (29)
Increasing the array size will reduce the quantization step of $\hat{\gamma}$, thereby improving polarization-state synthesis precision and enhancing XPL performance.
These can lead to the following conclusions. For a given polarization state and beam direction:
  • All PCRPA pattern performance, except for the dual-beam XPL, exhibits array-size dependence and shows progressive improvement with the array size.
  • The XPL performance of PCRPA dual-beam synthesis demonstrates negligible dependence on the array size.
  • In small-to-medium scale arrays, both the single-beam XPL and dual-beam PSL of the PCRPA demonstrate evident sensitivity to the array size, with rapid performance enhancement observed as the size expands. This scaling characteristic strongly suggests the proposed PCRPA architecture for large-scale PPA systems.
V Experimental Results and Discussion
V-A Experimental Results
To verify the effectiveness of the proposed architecture of the PCRPA and synthesis methods, an experiment is carried out using an X-band $8\times 8$-unit dual-polarized microstrip patch array antenna, shown in Fig. 16. The operating frequency is 9.30 GHz. The measurement results include active element pattern data for all ports, as shown in Fig. 16b, which means that the whole array can be utilized as a fully digital system.
Figure 16: An X-band $8\times 8$-unit dual-polarized microstrip patch array antenna test in an anechoic chamber. (a) Front view. (b) Back view.
Fig. 17 shows the CoP and XP patterns under the conventional and proposed architecture for single polarized pattern synthesis, with the beam direction at $(115^{\circ},10^{\circ})$ and the polarization state at V polarization. Compared with the CPPA, the PCRPA pattern mitigates the XPL from -19.67 dB to -37.75 dB, improving approximately 18 dB. Although this value indicates a degraded performance compared to the figure of simulated cases at -40 dB given in Fig. 15b, it still agrees with the previous conclusion of XPL suppression. Besides, the values of PSL for both conventional and proposed architectures remain comparable at -13.40 dB and -13.99 dB, respectively.
Figure 17: Measured CoP and XP patterns (elevation cut) under the conventional and proposed architecture for single polarized pattern synthesis.
Figure 18: Measured patterns (elevation cut) under different architectures for dual polarized pattern synthesis. (a) Conventional PPA. (b) Proposed PCRPA.
Fig. 18 shows the 2-D power patterns measured under two architectures for dual-polarized pattern synthesis, with the beam direction the same as the previous single-beam case. For the CoP patterns, both architectures demonstrate comparable matching performance: CPPA achieves a 3dB beamwidth of $5.79^{\circ}$ for H- and V-polarized beams, while PCRPA attains a narrower $4.98^{\circ}$ in both polarizations. The corresponding matching errors measure 5.35dB (CPPA) and 9.26dB (PCRPA). Compared to the CPPA, the CoP patterns of PCRPAs show PSL increase at 1.76 dB and directivity loss at 2.65 dB due to fewer channels. In addition, the XPL is effectively suppressed by calibration methods from -19.88 dB to -25.20 dB. The experiment results agree with the conclusions drawn by simulations but also reveal performance deterioration due to practical factors.
V-B Discussion
Analysis of both simulation and experimental data shows that the array size is a crucial factor for the performance of the proposed PCRPA architecture. Employing PCRPAs in larger arrays with a minimum size of $8\times 8$ is recommended. As the array size increases, the proposed PCRPAs exhibit lower sidelobe levels and better control over cross-polarization levels of the patterns. This can be attributed to several reasons:
  1. 1.
    The PCRPA architecture is subject to quantization errors, which decrease as the array size increases. This kind of error will limit the effects of XPL mitigation as the desired polarization state cannot be precisely synthesized. Larger arrays allow for finer resolution, thereby reducing the impact of quantization.
  2. 2.
    The pattern synthesis for PCRPAs is affected by various practical factors such as mutual coupling. The increase in the array size tends to mitigate these effects, leading to more precise and desirable pattern synthesis results.
  3. 3.
    For the dual polarized pattern synthesis, larger arrays can offer more DOFs in the thinned array synthesis problem, allowing for higher optimization limits on the PSL of the CoP patterns.
Furthermore, this architecture’s main advantage is its low cost compared to the conventional design. When the array size increases, the cost contradiction becomes more apparent. This also determines that the proposed PCRPA architecture is more suitable for addressing the cost-performance ratio issues in large-scale polarimetric phased array systems.
VI Conclusion
An architecture called PCRPA and its efficient pattern synthesis methods are proposed. Compared to the conventional architecture, the number of T/R channels in PCRPA is reduced by half using polarization coding reconfigurable elements. By simultaneously controlling the code and weight of each element, synthesis methods obtain arbitrarily or dual-polarized beam patterns with required beam directions, PSL, and XPL in an acceptable running time. This provides a low-cost scheme for PPA radars or other applications to save costs while ensuring beamforming abilities.
Despite the advantages, simulation results show that the proposed technique exhibits sensitivity to array size in terms of PSL and XPL in the pattern synthesis. Especially for XPL, the mitigation effect is influenced by multiple factors, such as quantization errors and mutual coupling. Moreover, due to the reduction in the number of T/R channels, there is an inevitable loss in beam power and directivity. Future research will focus on these problems.

### v3

@@
 Abstract
-Polarimetric phased arrays (PPAs) enhance radar target detection and anti-jamming capabilities.
-However, the dual transmit/receive (T/R) channel requirement leads to high costs and system complexity.
-To address this, this paper introduces a polarization-coding reconfigurable phased array (PCRPA) and associated pattern synthesis techniques to reduce PPA costs while minimizing performance degradation.
-Each PCRPA element connects to a single T/R channel and incorporates two-level RF switches for real-time control of polarization states and waveforms.
-By adjusting element codes and excitation weights, the PCRPA can generate arbitrarily polarized and dual-polarized beams.
-Efficient beam pattern synthesis methods are also proposed, featuring novel optimization constraints derived from theoretical and analytical analysis of PCRPAs.
-Simulations demonstrate that the approach achieves low cross-polarization and sidelobe levels comparable to conventional architectures within the scan range, particularly for large arrays.
-However, the channel reduction inevitably incurs power and directivity loss.
-Experiments conducted on an $8\times 8$ X-band array antenna validate the effectiveness of the proposed system.
-The PCRPA and synthesis methods are well-suited for large-scale PPA systems, offering significant cost-effectiveness while maintaining good sidelobe suppression and polarization control performance.
-Index Terms:
-Polarimetric phased array radar, low-cost phased array, polarization pattern synthesis, cross-polarization level (XPL).
+Polarimetric phased arrays (PPAs) enhance radar target detection and anti-jamming capabilities, but their conventional dual transmit/receive (T/R) channel architecture leads to high cost and system complexity.
+To address these limitations, this paper proposes a polarization-coding reconfigurable phased array (PCRPA) and associated pattern synthesis techniques, which reduce the channel count while preserving key performance.
+In the PCRPA, each antenna element connects to a single T/R channel and is equipped with a two-level RF switch, enabling real-time control of its polarization state and subarray grouping.
+By optimizing both the element polarization codes and the excitation weights, the array can synthesize arbitrarily polarized and dual-polarized beams.
+Simulation results show that the proposed approach achieves suppressed cross-polarization and comparable sidelobe levels compared to conventional PPAs across a wide scan range, with performance improvements being more pronounced in larger arrays.
+The inherent channel reduction does, however, incur a trade-off in terms of radiated power and directivity.
+Experimental validation using an $8\times 8$ X-band array antenna confirms the feasibility and effectiveness of the proposed system.
+The PCRPA architecture and the accompanying synthesis methods offer a cost-effective solution for large-scale PPA systems, maintaining sidelobe and polarization control with significantly reduced hardware complexity.
 publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@
 Polarimetric phased arrays (PPA) tend to be adopted by many radar systems to enhance their measurement performance within complex environments.
-The extra dimension of polarization can provide detailed information to detect targets in the presence of strong clutter[1, 2, 3, 4, 5], interference[6, 7, 8, 9, 10], or other adjacent objects[11, 12, 13].
+The extra dimension of polarization can provide detailed information to detect targets in the presence of strong clutter[hurtado2008polar, li2014ground, cheng2015antenna, wu2024a, an2024ground], interference[compton1981on, lu2021adaptive, yin2022radio, pan2023joint, an2024radio, xie2025coarray], or other adjacent objects[zhu2024principles, zhu2024a, liu2024robust].
 These applications rely on the beamforming capabilities of dual- or arbitrarily polarized patterns, which require independent channels of orthogonal polarization states, usually horizontal (H) and vertical (V).
-Therefore, conventional PPA architectures generally consist of dual-polarized antenna elements with transmit/receive (T/R) channels connected to each of their ports[14, 15, 16, 17].
+Therefore, conventional PPA architectures generally consist of dual-polarized antenna elements with transmit/receive (T/R) channels connected to each of their ports[salazar2015transmit, ta2015crossed, salazar2015tr, mirmozafari2019dual].
 Compared with single-polarized phased arrays of the same size, the number of channels in the T/R modules of PPAs is doubled.
-The conventional architectures are high-configuration but add to the burden of manufacturing the PPA systems, especially for large-scale arrays.
-A typical cost distribution for phased arrays given by [18] shows that the T/R modules can account for half the system cost, whose components grow proportionally with the number of elements.
-Meanwhile, many recently reported advanced PPA radar systems are on the scale of 1000-element or even bigger[19, 20, 21] to obtain high space resolution and more degrees of freedom (DOFs) for signal processing.
-Therefore, for these large-scale PPA systems, the heavy use of T/R channels under conventional architectures will make the total cost and complexity fairly high.
-And this also reveals the practical opportunity for cost reduction.
+Conventional architectures offer high performance but significantly increase the manufacturing cost and complexity of PPA systems, especially for large-scale arrays.
+A typical cost distribution for phased arrays given by [herd2016the] shows that the T/R modules can account for half the system cost, whose components grow proportionally with the number of elements.
+Meanwhile, many recently reported advanced PPA radar systems are on the scale of 1000-element or even bigger[fulton2017cylindrical, zhang2022cylindrical, palmer2023horus] to obtain high space resolution and more degrees of freedom (DoFs) for signal processing.
+Therefore, for these large-scale PPA systems, the heavy use of T/R channels under conventional architectures will make the total cost and complexity fairly high, which also reveals the practical opportunity for cost reduction.
 Over the past few decades, techniques like sparse arrays, thinned arrays, and subarrays have been thoroughly studied.
-They can significantly reduce the cost and complexity of phased arrays and directly apply to PPAs[22, 23].
+They can significantly reduce the cost and complexity of phased arrays and directly apply to PPAs[rocca2016uncon, mailloux2009irreg].
 However, these array designs are relatively fixed and static, so the configuration must be determined before manufacturing and cannot be changed.
 This will limit some crucial performance of radar systems, especially the power and scan range.
-In addition, polarization reconfigurable antennas are generally designed with single ports, requiring half the T/R channels compared to dual-polarized antennas[24, 25, 26].
+In addition, polarization reconfigurable antennas are generally designed with single ports, requiring half the T/R channels compared to dual-polarized antennas[nishamol2011an, lin2017multi, chen2021a].
 If they are employed as PPA elements, the system cost and complexity will also be significantly reduced.
@@
 Current research related to this issue can be divided into two schemes.
-The first has thoroughly investigated the application of polarization-reconfigurable elements in metasurfaces[27, 28], transmitarrays[29], and other non-electronically scanned array antennas[30, 31].
+The first has thoroughly investigated the application of polarization-reconfigurable elements in metasurfaces[yang2021recon, yang2023a], transmitarrays[wan2023code], and other non-electronically scanned array antennas[chen2022, wang2022].
 These works are instructive, but their array architectures and pattern synthesis models differ from the PPAs.
-The second scheme explores the idea of mixed polarized elements in PPA pattern synthesis but does not involve specific architecture design or beamforming requirements[32, 33, 34, 35, 36].
-For example, in [33, 34], a signal processing method called cross-polarization canceller (XPC) uses a selection of cross-polarized elements to improve the cross-polarization level (XPL) of beam patterns by canceling the corresponding field components.
+The second scheme explores the idea of mixed polarized elements in PPA pattern synthesis but does not involve specific architecture design or beamforming requirements[barbetty2012, salazar2024, salazar2024a, wang2024low-cost, zhou2024].
+For example, in [salazar2024, salazar2024a], a signal processing method called cross-polarization canceller (XPC) uses a selection of cross-polarized elements to improve the cross-polarization level (XPL) of beam patterns by canceling the corresponding field components.
 However, it is based on the conventional architecture and requires the phased array to be fully digital.
-Several architecture models have been reported in [35, 36], but they do not cover the commonly used modes of PPA radars, such as alternately transmitting and simultaneously receiving (ATSR) and simultaneously transmitting and simultaneously receiving (STSR).
+Several architecture models have been reported in [wang2024low-cost, zhou2024], but they do not cover the commonly used modes of PPA radars, such as alternately transmitting and simultaneously receiving (ATSR) and simultaneously transmitting and simultaneously receiving (STSR).
 This paper proposes an architecture and corresponding polarization pattern synthesis methods for PPA radars to reduce the cost and complexity while ensuring beamforming performance.
@@
 Given that the polarization states of each element are limited and independently controllable, this architecture is termed the polarization coding reconfigurable phased array (PCRPA).
-In the hardware design, each T/R channel has two switches: the antenna-connected switch for the antenna polarization state selection and the transceiver-connected switch for the waveform selection.
+In the hardware design, each T/R channel has two switches: the antenna-connected switch for the antenna polarization state selection and the transceiver-connected switch for the subarray selection.
 This paper proposes novel constraints, derived from antenna radiation and array geometry analysis, to optimize pattern performance and enhance algorithm efficiency.
@@
     A novel array architecture considering multiple applications for PPA radars is proposed.
-The designed architecture enables the real-time coding of polarization states and waveforms for reconfigurable antenna elements, allowing for the synthesis of arbitrarily polarized and dual-polarized beam patterns.
+The designed architecture enables the real-time coding of polarization states and subarrays for reconfigurable antenna elements, allowing for the synthesis of arbitrarily polarized and dual-polarized beam patterns.
 Compared to conventional architectures, the proposed design offers a cost-effective advantage by reducing the number of active channels in the T/R modules by half.
@@
 The remainder of this paper is organized as follows.
-Section II introduces the general design of the PCRPA and its beamformer architecture.
-Section III proposes the methods of polarization pattern synthesis based on the PRCA and verifies them through simulations.
-Section IV analyzes the parameter sensitivity of pattern synthesis, including the parameters of polarization state, array size, and beam position.
-Section V presents the experiment results to verify further the effectiveness of the proposed architecture and synthesis methods concerning PPA radars, then discusses the applicable scenarios.
-Section VI concludes the entire paper.
+Section II details the architecture design of the proposed PCRPA.
+Section III presents the theoretical foundation, formulating the efficient pattern synthesis methods.
+Section IV provides a comprehensive simulation validation and performance analysis, including parameter sensitivity studies.
+Section V offers experimental verification on a physical array prototype, discussing the agreement between measured and simulated results.
+Finally, Section VI concludes the paper.
 Notations: Some important symbols and notational definitions used in this article are given in Table I.
@@
 II General Architecture of The PCRPA
-Figure 1: The illustration of the conventional and the proposed polarimetric phased array architectures.
-(a) Architecture of the conventional polarimetric phased array, where two polarization ports of each element are connected to independent channels.
-(b) Architecture of the proposed polarimetric phased array with polarization reconfigurable elements.
-(c) Schematic diagram of the proposed array, which can produce dual-polarized or arbitrarily polarized beams by controlling the polarization states and waveforms of elements.
-The architecture designs of the conventional PPA and the proposed PCRPA are illustrated in Fig.
-1a and 1b.
-Although both diagrams adopt uniform $N$-element linear arrays with the spacing $d$ for simplicity, the proposed architecture can also be utilized in various antennas, including planar and conformal arrays.
-The elements for the conventional PPA are dual-polarized antennas consisting of two orthogonal polarization ports (usually H and V polarization), represented by the red and blue rectangles in Fig.
-1a.
-Regardless of the subarray design, each port is connected to an independent T/R channel, corresponding to the independent weighting $a_{ih}\phi_{ih}$ or $a_{iv}\phi_{iv}$ and waveform H or V.
-In this case, the electric field of beams formed by H and V ports can be expressed as:
+Figure 1: Illustration of the conventional and the proposed PPA architectures.
+(a) Conventional architecture.
+(b) Proposed architecture.
+(c) Schematic diagram of the proposed architecture, which can produce dual-polarized or arbitrarily polarized beams by controlling the polarization states and subarrays of elements.
+This section begins with an overview of the conventional PPA architecture.
+As illustrated in Fig.
+1a, the architecture consists of $N$ elements spaced at a uniform distance $d$ on the YOZ plane.
+In this architecture, the H and V polarization ports of each dual-polarized antenna element are connected to independent T/R channels (demonstrated by weight $a_{ih}\phi_{ih}$ or $a_{iv}\phi_{iv}$) and are fixedly assigned to corresponding H-polarized or V-polarized subarrays.
+The system synthesizes beams independently within the H and V subarrays via separate beamforming networks, which are then coherently combined in the far-field to achieve dual-polarized beam calibration or to synthesize beams with arbitrary polarization states.
+In this case, the electric field of beams formed by H and V subarrays can be expressed as:
 $\left\{{\begin{array}[]{l}{\boldsymbol{E}_{h}\left({\theta,\phi}\right)=\sum_{i=1}^{N}\boldsymbol{e}_{ih}\left({\theta,\phi}\right)}\\ {\boldsymbol{E}_{v}\left({\theta,\phi}\right)=\sum_{i=1}^{N}\boldsymbol{e}_{iv}\left({\theta,\phi}\right)}\end{array}}\right$.
@@
 where $\boldsymbol{e}_{ih}(\theta,\phi)$ and $\boldsymbol{e}_{iv}(\theta,\phi)$ denote the electric field vectors of the $i$-th H and V ports, and the effects of amplitude weights $a_{ih}$, $a_{iv}$ and phase weights $\phi_{ih}$, $\phi_{iv}$ have already been included.
-For the proposed PCRPA architecture, the elements are polarization reconfigurable and represented by the spliced rectangles in Fig.
-1b.
-The colored and transparent halves of the rectangles indicate polarization excited and unexcited, respectively.
-These elements are each connected to a single T/R channel, corresponding to the independent weighting $a_{i}\phi_{i}$ and waveform H or V.
-Compared to the conventional PPA, the number of T/R channels here is reduced to half.
-The PCRPA can form multiple polarized beams by controlling the coding states of elements, where the beam can be expressed as:
+However, this conventional approach requires a large number of TR channels, leading to high system cost and complexity.
+To address this issue, this paper proposes a PCRPA architecture, as illustrated in Fig.
+1b, where the number and positions of elements remain consistent with those in Fig.
+1a.
+The idea for cost reduction is achieved through the resource sharing of T/R channels among the polarization ports and subarray modules.
+Specifically, polarization-reconfigurable antenna elements are employed, allowing the dual-polarized ports of each element to time-share a single T/R channel, thereby halving the total number of channels.
+Furthermore, to address the beam reconfiguration requirements for dual-polarized pattern synthesis, a similar approach is applied, enabling the H- and V-polarized subarrays to also time-share a T/R channel.
+These two levels of reconfiguration collectively form an array architecture in which both the polarization and subarray connections are dynamically reconfigurable.
+In this way, the PCRPA architecture realizes beam synthesis through a dual-coding mechanism.
+First, subarray coding determines the set of element indices, denoted as $\boldsymbol{I}_{j}$, excited for the $j$-th beam.
+Second, polarization coding defines the sets of indices for all elements in the array excited through their H ports and V ports, denoted as $\boldsymbol{I}_{h}$ and $\boldsymbol{I}_{v}$.
+The total radiation pattern of beam $j$ is the superposition of the fields radiated by the elements in the intersection set of $\boldsymbol{I}_{jh}$ and $\boldsymbol{I}_{jv}$, defined as:
+$\left\{{\begin{array}[]{l}{\boldsymbol{I}_{jh}=\boldsymbol{I}_{j}\cap\boldsymbol{I}_{h}}\\ {\boldsymbol{I}_{jv}=\boldsymbol{I}_{j}\cap\boldsymbol{I}_{v}}\end{array}}\right$. (2)
+And the electric field of the $j$-th beam can be expressed as:
 $\boldsymbol{E}_{j}(\theta,\phi)=\sum_{i=1}^{N_{jh}}\boldsymbol{e}_{I^{jh}_{i}h}(\theta,\phi)+\sum_{i=1}^{N_{jv}}\boldsymbol{e}_{I^{jv}_{i}v}(\theta,\phi)$ (3)
-where $\boldsymbol{E}_{j}(\theta,\phi)$ is the electric field of the $j$-th beam, $\boldsymbol{I}^{jh}$ and $\boldsymbol{I}^{jv}$ are vectors containing the element indices coded by H and V polarization for the $j$-th beam, $N_{jh}$ and $N_{jv}$ are the number of these elements.
+where $N_{jh}$ and $N_{jv}$ are the number of elements in sets $\boldsymbol{I}_{jh}$ and $\boldsymbol{I}_{jv}$.
 Here, vectors $\boldsymbol{I}^{jh}$ and $\boldsymbol{I}^{jv}$ should satisfy the constraints followed:
-$\left\{{\begin{array}[]{l}{1\leq I^{jA}_{i}\leq N,\>I^{jA}_{i}\in\mathbb{Z},i=1,2,\cdots,N_{jA}}\\ {I^{jA}_{m}\neq I^{jA}_{n},\>\forall m,n\in\{1,2,\cdots,N_{jA}\},m\neq n}\\ {\boldsymbol{I^{jh}\cap\boldsymbol{I}^{jv}=\emptyset}}\end{array}}\right$.
+$\left\{{\begin{array}[]{l}{1\leq I^{jA}_{i}\leq N,\>I^{jA}_{i}\in\mathbb{Z},i=1,2,\cdots,N_{jA}}\\ {I^{jA}_{m}\neq I^{jA}_{n},\>\forall m,n\in\{1,2,\cdots,N_{jA}\},m\neq n}\\ {\boldsymbol{I}^{jh}\cap\boldsymbol{I}^{jv}=\varnothing}\end{array}}\right$.
 (4)
 where $A\in\{h,v\}$ represents the polarization port.
-Therefore, it can be summarized from Fig.
-1b and Equation (2)-(3) that beam patterns of the PCRPA are synthesized by 3 groups of variable: the amplitude and phase weights, the waveform code and the polarization code.
-The latter two groups exist only in the proposed architecture rather than the conventional PPA.
-By adjusting these variables in real-time, the PCRPA can form multiple electronically scanned beams with any desired polarization.
-Considering the practical scenarios in radar applications, this paper focuses on two problems: arbitrarily polarized and dual-polarized pattern synthesis, shown in Fig.
+Based on Equation 3, the proposed PCRPA architecture is able to form electrially scanned beams of either arbitrary or dual polarization, shown in Fig.
 1c.
-These beam patterns of high performance can meet the requirements of polarization measurements and anti-interference.
-Since the proposed PCRPA architecture includes a two-level coding process, it can be realized by the beamformer designed with two-level RF switches shown in Fig.
-2.
-Here, RF switches are utilized to connect the array antenna, T/R modules, and transceivers sequentially.
-To highlight the architectural features, the radar transmitters and receivers are schematically represented in the diagram.
-Regarding digitization levels: The current design reflects the active phased array architecture; Subarray digital architecture can be achieved by incorporating multiple digital transceivers after the T/R modules; For fully digital architectures, the second-level switches become redundant and can be eliminated[37].
-This is because the waveforms can be conveniently altered by direct digital synthesizers (DDS) in digital transceivers at the element level.
-Besides, the combination of dual-polarized antennas and RF switches can also be replaced by polarization reconfigurable antennas with single ports.
-However, the latter may not tolerate high transmitting power due to components like PIN diodes.
-Figure 2: The beamformer design of the proposed phased array architecture for active arrays or digitized subarrays, where RF switches are utilized between T/R modules, array antennas, and transceivers.
+This is owing to two key aspects: Firstly, the code vectors $\boldsymbol{I}_{jh}$ and $\boldsymbol{I}_{jv}$ create irregular distributions of phase centers across the aperture, analogous to the synthesis principles of sparse arrays, which enables the formation of high-gain, low-sidelobe beams.
+Secondly, by dynamically configuring the number and excitation coefficients (amplitude and phase) of elements fed through different polarization ports, the architecture can achieve functionality similar to cross-polarization calibration techniques, allowing for the precise synthesis of any desired polarization state.
+Figure 2: The beamformer design of the proposed phased array architecture for different phased array architectures. (a) For subarray-level digital phased array. (b) For fully digital phased array.
+TABLE II: Comparison of PPA Architectures
+Architecture Channel Number Element Polarization Beamforming DoFs Arbitrarily Polarized Beamforming Dual Polarized Beamforming
+Conventional $2N$ Fixed Weight Yes Yes
+Ref. [zhou2024] NN Fixed Weight Yes No
+Proposed NN Reconfigurable Weight, Port, Subarray Yes Yes
+In terms of the hardware implementation, the two-level coding process of the PCRPA can be realized by the two-level of RF switches shown in Fig.
+2a.
+Here, RF switches are utilized to connect the array antenna, T/R modules, and beamformers sequentially.
+Here, the combination of dual-polarized antennas and RF switches may also be replaced by polarization reconfigurable antennas with single ports.
+However, the latter may not tolerate high transmitting power of radar systems at presence due to components like PIN diodes.
+Besides, although this paper focuses on the subarray-level digital systems, the PCRPA is also applicable to fully digital phased arrays.
+In a fully digital architecture shown by Fig 2b, the subarray coding function can be implemented directly by the digital beamformer[fulton2016digital], allowing the second-level RF switch to be omitted, thereby simplifying the hardware structure.
+To summarize, a comparative analysis is provided by Table II.
+In comparison to the conventional PPAs, both the proposed PCRPA architecture and a previously reported structure by the team [zhou2024] can reduce the number of T/R channels by half, lowering the system cost and complexity.
+Relative to the prior work, the key difference of the PCRPA lies in beamforming DoFs and capabilities.
+While [zhou2024] utilized fixed phase shifters to decide element polarization, this work adopted a real-time reconfigurable design.
+This makes the port polarization a new DoF for beamforming in the proposed architecture.
+Additionally, the subarray coding DoF is introduced.
+These enhancements enable the PCRPA to achieve dual-polarized beamforming.
+Meanwhile, the real-time reconfigurability of polarization coding also leads to better radiation performance, which will be demonstrated later.
 III Pattern Synthesis for The PCRPA
@@
 where $A,B\in\{h,v\}$ represent the polarization of the electric field component and the antenna port, respectively.
-The beam patterns are formed for the active phased arrays by adding each AEP with independent weights. In this case, the array pattern can be expressed as[38]:
+The beam patterns are formed for the active phased arrays by adding each AEP with independent weights. In this case, the array pattern can be expressed as[pozar1994the]:
 $\left\{{\begin{array}[]{l}{{f_{h}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{g}_{hh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{g}_{hv}(\theta,\phi)}\\ {{f_{v}}\left({\theta,\phi}\right)=\boldsymbol{w}_{h}^{T}\boldsymbol{g}_{vh}(\theta,\phi)+\boldsymbol{w}_{v}^{T}\boldsymbol{g}_{vv}(\theta,\phi)}\end{array}}\right$.
@@
 Figure 3: Flowchart of the arbitrarily polarized pattern synthesis method.
-For arbitrarily polarized pattern synthesis problems, define the desired polarization basis according to Ludwig’s polarization definition II[39]:
+For arbitrarily polarized pattern synthesis problems, define the desired polarization basis according to Ludwig’s polarization definition II[ludwig1973the]:
 $\left\{\begin{array}[]{l}{\boldsymbol{e}_{co}=\left[\cos{\gamma,\sin{\gamma}e^{j\eta}}\right]^{T}}\\ {\boldsymbol{e}_{cr}=\left[-\sin{\gamma,\cos{\gamma}e^{-j\eta}}\right]^{T}}\\ {\boldsymbol{e}_{co}^{H}\boldsymbol{e}_{cr}=0}\end{array}\right$.
@@
 $=\left[\begin{array}[]{l}{\cos(\gamma)}\\ {\sin(\gamma)e^{j\eta}}\end{array}\right]$
-Based on Equation (13), the conventional PPAs can synthesize patterns of arbitrarily polarized beams by directly multiplying the corresponding weights $\cos{\gamma}$ and $\sin{\gamma}e^{-j\eta}$ into the polarization channels H and V[36].
+Based on Equation (14), the conventional PPAs can synthesize patterns of arbitrarily polarized beams by directly multiplying the corresponding weights $\cos{\gamma}$ and $\sin{\gamma}e^{-j\eta}$ into the polarization channels H and V[zhou2024].
 However, this method ignores the cross-polarization components radiated at the H and V ports.
@@
 Figure 4: Histogram of the Monte Carlo simulation results concerning arbitrarily polarized pattern synthesis.
-The whole method can be summarized step by step, given in Algorithm 1.
-Although Equation (18) can uniquely calculate the number of H and V elements, this method provides an adjustment mechanism realized by Steps 4 and 7.
-This is because the AEPs of embedded elements may not fully represent the array polarization characteristics due to mutual coupling and other effects.
-For each $n_{h}$ and $n_{v}$, a maximum generation is set.
-If the iteration reaches the maximum generation but the patterns do not meet the requirements, the method will adjust $n_{h}$ and $n_{v}$.
-The adjustment rules can be flexible, for example, reducing one H element each time.
-The random number generator in Step 5 is chosen according to simulation results given by Fig.
-4, which can also be replaced by other optimization algorithms if needed.
+The proposed algorithm, outlined in Algorithm 1, takes the AEPs of all ports, along with performance specifications (e.g., beam steering, polarization, and sidelobe level), and outputs the polarization code and weight vectors accroding to Fig.
+3.
+A central challenge addressed by this method is the mitigation of mutual coupling effects between elements and ports.
+To this end, the algorithm incorporates a two-fold strategy: First, while the polarization basis in Step 2 is constructed using the data of the center embedded element, the array pattern synthesis in Step 5 is computed based on the full AEP model.
+This model utilizes the full-wave simulated or measured data of all ports, thereby accounting for mutual coupling.
+Second, to compensate for any residual inaccuracies arising from the single-element estimation in Step 2, Step 7 provides an adjustment mechanism.
+Specifically, if the pattern requirements are not met upon reaching the maximum generation for a given $n_{h}$ and $n_{v}$, the ratio of H-to-V elements is adaptively modified according to a predefined rule (e.g., reducing one H element per adjustment).
+Furthermore, the random number generator in Step 5, whose distribution is informed by the Monte Carlo simulation results in Fig.
+4, can be seamlessly replaced by other optimization algorithms for initial value generation if required.
 Algorithm 1 The proposed arbitrarily polarized pattern synthesis method for the PCRPA.
@@
 Therefore, all the elements are split into two groups to synthesize the beams.
-This can be represented by the waveform coding vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, where $x_{1},x_{2}\in\{0,1\}$ and $x_{i,j}=1$ means the $i$-th element belongs to the $j$-th beam.
+This can be represented by the subarray coding vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, where $x_{1},x_{2}\in\{0,1\}$ and $x_{i,j}=1$ means the $i$-th element belongs to the $j$-th beam.
 Take the beam $1$ as an example, Equation (11) can be transferred to:
@@
 (23)
-In this case, the synthesis of dual-polarized beam patterns is decided by variables of the beam coding vectors $\boldsymbol{x}_{1h}$, $\boldsymbol{x}_{1v}$, $\boldsymbol{x}_{2h}$, and $\boldsymbol{x}_{2v}$, which are defined by the Hadamard product of the polarization and waveform coding vectors.
-According to [40], the calibration methods can conveniently mitigate the cross-polarization components of these dual-polarized beams.
-Therefore, it is reasonable to assume that no cross-polarized elements are involved in each beam, which can be expressed as:
-$\left\{{\begin{array}[]{l}{\boldsymbol{x}_{1}=\boldsymbol{x}_{h}}\\ {\boldsymbol{x}_{2}=\boldsymbol{x}_{v}}\end{array}}\right$. (23)
-Since Equation (23) holds, the variables $\boldsymbol{x}_{1v}$ and $\boldsymbol{x}_{2h}$ are equal to zero, from which can infer that:
-$\boldsymbol{x}_{1h}+\boldsymbol{x}_{2v}=\boldsymbol{1}$ (24)
-where $\text{sum}(\boldsymbol{x}_{1h})=\text{sum}(\boldsymbol{x}_{2v})=N/2$ to ensure good match of the co-polar beams.
+The model given in Equation (23) involves nonlinear vector dot products, making direct solution computationally complex.
+A divide-and-conquer approach is therefore adopted.
+The problem is decomposed into two sequential steps: first, the subarray coding vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{1}$, which determine the co-polar pattern performance for dual beams, are synthesized.
+Subsequently, the polarization coding vectors $\boldsymbol{x}_{h}$ and $\boldsymbol{x}_{v}$ are optimized using Algorithm 1 described in the previous subsection for arbitrary polarization pattern synthesis, with detailed steps provided later.
 A Monte Carlo simulation given by Fig.
-5 has been carried out according to Equation (24) to investigate the PSL distribution under the dual-polarized pattern synthesis model.
+5 has been carried out to investigate the PSL distribution under the dual-polarized pattern synthesis model.
+According to the divide-and-conquer approach, assume that $\boldsymbol{x}_{1}=\boldsymbol{x}_{h}=\boldsymbol{x}_{1h}$ and $\boldsymbol{x}_{2}=\boldsymbol{x}_{v}=\boldsymbol{x}_{2v}$.
 The beampoint is set at $(120^{\circ},15^{\circ})$, with the PSL taking the maximum of the H and V beams.
@@
 Therefore, an optimized model with a central symmetry constraint is introduced to derive low PSL results.
-This model can be solved using binary genetic algorithm (BGA) methods and has been discussed in detail in previous work by the authors[35].
+This model can be solved using binary genetic algorithm (BGA) methods and has been discussed in detail in previous work by the authors[wang2024low-cost].
 The central symmetry constraint means that two configurations forming the H and V beams are centrally symmetric shown in Fig.
 6.
-In this case, the array factors produced by these configurations will have the same power patterns, which have been theoretically demonstrated in [35].
+In this case, the array factors produced by these configurations will have the same power patterns, which have been theoretically demonstrated in [wang2024low-cost].
 Figure 6: Schematic diagram of the central symmetry constraint.
 This property ensures the optimization model derives highly matched co-polar beam patterns. The constraint can be expressed by the following equation:
-$\boldsymbol{X}_{2v}=\boldsymbol{1}-\text{rot180}(\boldsymbol{X}_{1h})$ (25)
+$\boldsymbol{X}_{2}=\boldsymbol{1}-\text{rot180}(\boldsymbol{X}_{1})$ (24)
 where $\text{rot180(}\cdot\text{)}$ means to rotate a matrix 180 degrees.
-From the perspective of matrix operations, Equation (25) can be further represented in vector form using the code vectors $\boldsymbol{x}_{1h}$ and $\boldsymbol{x}_{2v}$:
+From the perspective of matrix operations, Equation (24) can be further represented in vector form using the subarray code vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$:
 […]
@@
 In this paper, the 1-elements are arranged on the subdiagonal of the matrix.
-The pattern synthesis problem can be then formulated as a minimax problem for minimizing the PSL of $f_{1h}\left(\theta,\phi\right)$ and $f_{2v}\left(\theta,\phi\right)$ by determining the coding of elements in the binary vector $\boldsymbol{x}_{u}$, which is given by
+The pattern synthesis problem can be then formulated as a minimax problem for minimizing the PSL of $f_{1}\left(\theta,\phi\right)$ and $f_{2}\left(\theta,\phi\right)$ by determining the coding of elements in the binary vector $\boldsymbol{x}_{u}$, which is given by
 […]
-This nonlinear integer programming problem tends to be solved by stochastic optimization algorithms such as BGA, binary PSO (BPSO), and binary differential evolution (BDE).
-In this paper, the BGA algorithm is adopted.
-The whole procedure of the synthesis method is shown in Algorithm 2 and explained step by step.
+This nonlinear integer programming problem is solved using a stochastic optimization algorithm, specifically the BGA adopted in this work.
+For the $i$-th subarray coding vector $\mathbf{x}_{i}\in\{0,1\}^{n}$, a transformation matrix $\mathbf{L}_{i}\in\{0,1\}^{n_{i}\times n}$ is constructed, where $n_{i}=\|\mathbf{x}_{i}\|_{1}$ denotes the number of active elements in this subarray.
+Let $\mathbf{Q}=\{q_{1},q_{2},\ldots,q_{n_{i}}\}$ be the ordered set of indices where the components of $\mathbf{x}_{i}$ are equal to 1.
+The transformation matrix $\mathbf{L}_{i}$ is then defined element-wise by:
+[…]
+for row index $k=1,2,\ldots,n_{i}$ and column index $l=1,2,\ldots,n$. This matrix maps from the full array’s indices to the indices of the given subarray code vector.
+Since the dual-beam synthesis problem is decomposed into two independent single-beam synthesis subproblems, for the $i$-th beam ($i=1,2$), the input data can be obtained by mapping the original AEP data $\mathbf{g}_{AB}$ via the transformation matrix $\mathbf{L}_{i}$:
+$\tilde{\mathbf{g}}_{AB}^{(i)}(\theta,\phi)=\mathbf{L}_{i}\mathbf{g}_{AB}(\theta,\phi)$ (28)
+Given the constraints including the beam direction, polarization state, and sidelobe level, Algorithm 1 is performed to optimize $\tilde{\mathbf{g}}_{AB}^{(i)}(\theta,\phi)$, yielding the local solutions $\mathbf{x}_{h}^{i}$ and $\mathbf{x}_{v}^{i}$ corresponding to this subarray.
+Upon completion of the optimization for all beams, the local solutions are combined into the final code vectors for the full array via the transpose of the transformation matrices:
+$\left\{\begin{aligned} \mathbf{x}_{ih}=\mathbf{L}_{i}^{\text{T}}\mathbf{x}_{h}^{i}\\ \mathbf{x}_{iv}=\mathbf{L}_{i}^{\text{T}}\mathbf{x}_{v}^{i}\end{aligned}\right$. (29)
+The vectors $\mathbf{x}_{ih}$ and $\mathbf{x}_{iv}$ constitute the final output of the dual-beam pattern synthesis process.
+Consistent with the approach in Algorithm 1, the synthesis procedure given by Algorithm 2 is also founded on the AEP model.
+By utilizing full-wave simulation or measured data for array pattern calculation, this method inherently accounts for mutual coupling effects.
+The algorithm steps are explained as follows.
 Algorithm 2 The proposed dual-polarized pattern synthesis method for the PCRPA.
@@
 Set the termination condition of the run time or the maximum number of generations.
-2: Optimize the variable $\boldsymbol{x}_{u}$ using the BGA algorithm according to the model given by Equation (27). Stop the process if it reaches the termination condition.
-3: Calculate the optimized configuration results according to Equation (26).
-4: Calibrate the dual-polarized patterns according to the method mentioned in [40].
-III-D Simulation Results
-Input data from the HFSS model of a $16\times 16$-unit dual-polarized microstrip array antenna is adopted to demonstrate the performance of the proposed architecture and methods.
-The simulations are carried out for vertically, arbitrarily, and dual-polarized patterns concerning various radar applications.
-III-D1 Vertically Polarized Pattern
-This simulation considers a V polarization pattern with the beam direction at $(120^{\circ},15^{\circ})$.
-These beams with pure V or H polarization are standard in polarization-measuring applications of ATSR modes.
-According to the initial conditions and the method given in Algorithm 1, the results of the element polarization configuration are shown in Fig.
-7.
-In this configuration, 17 out of 256 elements, which account for 6.64%, are excited by the H-pol ports to eliminate the cross-polarization components of the synthesized beam.
-Their positions are generated based on the proposed method.
+2: Optimize the variable $\boldsymbol{x}_{u}$ using the BGA algorithm according to the model given by Equation (26).
+Stop the process if it reaches the termination condition and obtain the subarray code vectors $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$.
+3: Perform Algorithm 1 under $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ respectively according to Equation (27)-(29). Return the synthesis results.
+IV Simulation And Analysis
+This section presents simulation examples to validate the proposed PCRPA architecture and its pattern synthesis methods.
+The simulation begins with acquiring the AEP data of all ports from a dual-polarized array model through full-wave analysis using the high-frequency structure simulator (HFSS).
+A dual-polarized microstrip patch antenna, operating at a center frequency of 10 GHz, was employed as the array element.
+The square patch has a side length of 8.3 mm and is fed by coaxial probes.
+The array lies in the yoz-plane, with all elements arranged in a $16\times 16$ uniform rectangular lattice with an element spacing of 15 mm (i.e., half-wavelength at 10 GHz).
+Since this section focuses on validating the proposed architecture and methodology, the electromagnetic simulation was configured for a single frequency of 10 GHz.
+These data are then used as input to the synthesis model, where the binary coding vectors are optimized to emulate the switching functionalities of the architecture in Fig.
+2.
+A set of three key metrics is employed to evaluate the pattern performance: PSL, XPL, and directivity (Dir).
+PSL and Dir characterize the concentration of co-polarized radiation energy, while XPL quantifies the polarization purity at the beam direction.
+These metrics are consequently adopted for clear comparison in following simulations and experiments.
+IV-A Arbitrarily Polarized Pattern
+This set of simulations evaluates the performance of Algorithm 1 for synthesizing arbitrarily-polarized beams steered to $(120^{\circ},15^{\circ})$.
+The synthesis results of array configurations and beam patterns for two polarization states, $(90^{\circ},0^{\circ})$ and $(45^{\circ},0^{\circ})$, are presented in Fig.
+7 as examples.
+Figure 7: Simulation results for arbitrarily polarized pattern synthesis.
+(a) Array configuration (V-polarization).
+(b) Array configuration ($45^{\circ}$ linear polarization).
+(c) CoP pattern of V-polarized beam.
+(d) XP pattern of V-polarized beam.
+(e) CoP pattern of $45^{\circ}$ linearly-polarized beam.
+(f) XP pattern of $45^{\circ}$ linearly-polarized beam.
+In Fig.
+7a and 7b, orange and blue colors correspond to the H and V port excitations, respectively.
+The results show that the PCRPA uses 17 H ports and 239 V ports to synthesize the V-polarized beam, whereas for the $45^{\circ}$ linearly polarized beam, the numbers change to 159 and 97.
+Both configurations form stable main lobes at the desired direction with controlled PSL and obvious XPL suppression.
+However, the XP pattern of the $45^{\circ}$ linearly polarized beam exhibits a higher level across the entire space, indicating a sensitivity of the synthesis performance to the polarization state, which will be further discussed later.
 Fig.
-8 presents the 3-D normalized power patterns of different array architectures or methods under the same data input and initial conditions.
-The calibration method aimed at the ATSR mode is adopted from [40].
-Compared to the CPPA, PCRPA exhibits a noticeable improvement in XP component mitigation, with the value of the cross-polarization level XPL decreasing from -22.93 dB to -59.68 dB, as shown in Fig.
-8b and 8f.
-This is comparable to the effects of calibration, as shown in Fig.
-8d, where the XPL is -65.31 dB.
-Meanwhile, the CoP pattern of PCRPA indicates only an even better level in the PSL compared to conventional PPA, as shown in Fig.
-8a, 8c, and 8e.
-Figure 7: H-pol and V-pol element distribution of a $16\times 16$-unit planar array for the required polarization state $(90^{\circ},0^{\circ})$ and beam direction $(120^{\circ},15^{\circ})$.
-The blue and red squares represent H-pol and V-pol elements, respectively.
-Figure 8: The 3-D normalized power patterns of the simulated $16\times 16$-unit microstrip antenna array, where V polarization is the co-polarization and the beam direction is $(120^{\circ},15^{\circ})$.
-The red dashed line helps indicate the pointing direction of the array beam.
-(a) The CoP pattern before calibration of the CPPA.
-(b) The XP pattern before calibration of the CPPA.
-(c) The CoP pattern after calibration of the CPPA.
-(d) The XP pattern after calibration of the CPPA.
-(e) The CoP pattern of the proposed PCRPA.
-(f) The XP pattern of the proposed PCRPA.
-TABLE II: Comparison of the key parameters of the vertically polarized pattern synthesis
-Group CPPA, uncalibrated CPPA, calibrated PCRPA
-XPL/dB -22.93 -65.31 -59.68
-PSL/dB -12.75 -12.77 -13.39
-Pt/dB 0 -0.01 -0.55
-Dir/dBi 28.37 28.37 28.15
-To clarify in detail, Table III presents the specific parameter values of the three patterns.
-Here, $P_{t}$ represents the transmission power at the beam direction normalized with the value of the uncalibrated beam under CPPA; $D$ denotes the directivity of the array pattern.
-As indicated by the data, since some elements are synthesized to suppress XP components, the integrating sidelobes of both the CoP and XP patterns of PCRPA have risen, leading to a loss of power and directivity.
-The results show that $P_{t}$ and $D$ of PCRPA are reduced by 0.55 dB and 0.22 dB, respectively.
-III-D2 Arbitrarily Polarized Pattern
-This simulation considers an arbitrarily polarized pattern with the same beam direction as in the above section.
-Such requirements tend to appear in many anti-interference applications, where the beam polarization is orthogonal to the interference signal polarization to suppress its received power.
-Here, the desired polarization state is set to $(30^{\circ},60^{\circ})$.
-According to Equation (13), the 2-D normalized power patterns of the CPPA are given by Fig.
+8 further compares the elevation plane cuts of the PCRPA patterns from Fig.
+7 with those from the CPPA and [zhou2024], with the corresponding quantitative metrics summarized in TABLE III.
+The comparative results show that for V-polarized synthesis, both [zhou2024] and the PCRPA maintain a PSL comparable to the CPPA, albeit with directivity loss at approximately 0.2 dB due to elevated integrated sidelobes.
+In terms of cross-polarization, the architecture in [zhou2024] exhibits an XPL similar to the CPPA, with its XP pattern generally elevated across the entire spatial domain.
+In contrast, the PCRPA achieves significant suppression of both the XPL and the cross-polarization within the main lobe.
+For $45^{\circ}$ linearly-polarized synthesis, the PCRPA maintains the same advantages observed in the V-polarized case; however, its level of XP pattern outside the main lobe is further increased.
+It is noteworthy that the architecture proposed in [zhou2024] cannot form polarization states other than H and V under uniform weighting; therefore, its performance for $45^{\circ}$ linear polarization could not be compared under identical conditions.
+Figure 8: Normalized 2-D power patterns (elevation plane cut) for dual-polarized beams.
+(a) CoP patterns of V-polarized beams.
+(b) XP patterns of V-polarized beams.
+(c) CoP patterns of $45^{\circ}$ linearly-polarized beams.
+(d) XP patterns of $45^{\circ}$ linearly-polarized beams.
+TABLE III: Comparison of key metrics for the arbitrarily polarized pattern synthesis
+CPPA Ref. [zhou2024] PCRPA
+$(90^{\circ},0^{\circ})$ XPL (dB) -22.93 -22.89 -59.68
+PSL (dB) -12.75 -12.70 -13.39
+Dir (dBi) 28.37 28.14 28.15
+$(45^{\circ},0^{\circ})$ XPL (dB) -24.43 -49.13
+PSL (dB) -12.75 -13.38
+Dir (dBi) 28.37 28.00
+Amplitude Weight Uniform Uniform Uniform
+IV-B Dual-polarized Patterns
+This set of simulations evaluates the performance of the proposed dual-polarized beam synthesis method with beams steered to $(120^{\circ},15^{\circ})$.
+The sythesis results obtained are presented in Fig.
 9.
-The XPL of this beam is at a relatively high level of -21.25 dB, which may affect the anti-interference processing.
-Figure 9: The azimuth cut plane of the 2-D normalized power pattern of the arbitrarily polarized pattern synthesis under the CPPA architecture, with the array size of $16\times 16$, co-polarization state at $(30^{\circ},60^{\circ})$ and the beam direction at $(120^{\circ},15^{\circ})$.
+Diamonds and squares denote the subarray excitations for the H- and V-polarized beams, respectively; orange and blue colors correspond to the H and V port excitations of the elements.
+As shown, each beam is synthesized using half elements of the total array.
+The H-polarized beam is excited by 120 H ports and 8 V ports, whereas the V-polarized beam is excited by 120 V ports and 8 H ports.
+Fig 9 also presents the CoP and XP patterns for both beams.
+It can be observed that both CoP patterns form stable main lobes at the desired beam direction, with the PSL below -13 dB, and the XPL suppressed to a low level.
+Conversely, the integrated sidelobe levels of the CoP patterns are higher than those in the single-beam case, leading to a measurable loss in directivity.
+Figure 9: Simulation results for dual-polarized beam synthesis.
+(a) Array configuration.
+(b) CoP pattern of H-polarized beam.
+(c) XP pattern of H-polarized beam.
+(d) CoP pattern of the V-polarized beam.
+(e) XP pattern of the V-polarized beam.
 Fig.
-10 presents three groups of PCRPA results obtained by different pattern synthesis methods, including the proposed, random, and BGA methods.
-Table III lists the key parameters of these results and the CPPA in detail.
-Here, $\text{XPL}_{\text{m}}$ denotes the normalized peak level of the XP pattern across the scan principal planes.
-The difference between these three methods lies in Step 4 to Step 6 of the algorithm given in subsection III-B.
-The random method obtains the array configuration by performing the random number generator only once, where the XPL and PSL constraints are ineffective.
-The BGA method obtains the array configuration by optimizing the coding vectors $\boldsymbol{x}_{h}$ and $\boldsymbol{x}_{v}$ using a binary genetic algorithm.
-Computations were performed on a PC with an Intel Core i7-10510U central processing unit (CPU) at 2.30 GHz and 16 GB RAM.
-Figure 10: Comparison of the arbitrarily polarized pattern synthesis results for the proposed architecture using different methods.
-(a) Elevation plane using the proposed method.
-(b) Elevation plane using the random method.
-(c) Elevation plane using the GA method.
-(d) Azimuth plane using the proposed method.
-(e) Azimuth plane using the random method.
-(f) Azimuth plane using the GA method.
-(g) Array configuration using the proposed method.
-(h) Array configuration using the random method.
-(i) Array configuration using the GA method.
-TABLE III: Comparison of key parameters of the arbitrarily polarized pattern synthesis
-Group CPPA PCRPA
-proposed random BGA
-XPL/dB -24.43 -49.13 -49.13 -49.05
-$\text{XPL}_{\text{m}}$/dB -24.15 -20.29 -8.51 -20.61
-PSL/dB -12.75 -13.38 -12.63 -13.93
-Run time/s \ 5.19 \ 1206.95
-Considering the comparison of XPL and PSL between CPPA and PCRPA, the simulation results here are consistent with the V polarization simulations.
-The advantage of PCRPA’s cross-polarization suppression still exists, and so does the decrease of the peak sidelobe, which is at a value of -13.38 dB.
-Besides, among the 3 groups of PCRPA results, the XP patterns of the random method show an excessively high level at -8.71 dB within the main lobe area, given in Fig.
-10b and 10e.
-This excessively high sidelobe of XP patterns is eliminated in the other 2 groups due to the optimization of element configurations, as shown in Fig.
-10a, 10c, 10d, and 10f.
-Furthermore, the proposed method and the BGA method’s pattern synthesis results indicate comparable performance, with the proposed method’s PSL being 0.55 dB higher.
-However, its run time is also significantly shorter, only 5.19 seconds compared with 1206.95 seconds using the BGA method according to Table III, which is an obvious efficiency advantage.
-The simulation results presented above show that the proposed method can achieve a compromise between performance and efficiency in the PCRPA architecture.
-III-D3 Dual-polarized Patterns
-Figure 11: H-pol and V-pol element distribution of a $16\times 16$-unit planar array for dual-polarized beams with beam direction $(120^{\circ},15^{\circ})$.
-The blue squares and red diamonds represent H-pol and V-pol elements, respectively.
-Figure 12: Comparison of the dual-polarized pattern synthesis results for the conventional and proposed architecture using different methods.
-(a) Elevation plane of the CPPA.
-(b) Elevation plane of the PCRPA using the proposed method.
-(c) Elevation plane of the PCRPA using the random method.
-(d) Azimuth plane of the CPPA.
-(e) Azimuth plane of the PCRPA using the proposed method.
-(f) Azimuth plane of the PCRPA using the random method.
-TABLE IV: Comparison of different groups
-Group CPPA PCRPA
-proposed random
-XPL/dB H -56.89 -56.76 -56.86
-V -57.80 -57.87 -57.84
-PSL/dB H -12.74 -12.73 -11.70
-V -12.77 -12.76 -8.95
-$P_{t}$/dB H 0 -6.02 -6.02
-V -0.02 -6.04 -6.04
-DD/dBi H 28.37 26.12 25.86
-V 28.37 26.12 26.07
-ME/dB -19.57 -19.59 10.36
-This simulation considers a pair of H and V polarization beams with the same beam direction as above, which are commonly required in polarization measuring applications of the STSR mode.
-According to the proposed synthesis method given by Equation (27) and Algorithm 2, the entire array is split into two groups, each fed with the corresponding waveform.
-To quantitatively evaluate the matching performance of the main lobes in dual-polarized beams, a new parameter called matching error (ME) is introduced, defined as[41]:
-[…]
-where $f_{1h,k}$ and $f_{2v_{k}}$ represent the $k$-th sampling points within the main lobe of patterns $f_{1h}(\theta,\phi)$ and $f_{2v}(\theta,\phi)$.
-Fig.
-12 compares three groups of dual polarization patterns obtained under different configurations: the CPPA, the optimized PCRPA, and a bad case of PCRPA.
-All the patterns have been processed by polarization calibration methods.
-Fig.
-11 shows the configuration of the optimized PCRPA.
-Markers of squares and diamonds represent the H and V polarization excited elements to distinguish the coding of waveforms.
-The bad case configuration is obtained without the central symmetry constraint proposed in this paper, simply using a random number generator instead.
-The optimized PCRPA performs well on the matched co-polar beams shown in Fig.
-12b and 12e.
-Its ME value is very close to that of conventional PPA and 29.96 dB better than that of the bad case, which verifies the effectiveness of the constraint.
-The same phenomenon also applies to the PSL parameter detailed in Table IV.
-In addition to the above advantages, Table IV shows that the PCRPA has a significant loss problem on power and directivity at around 6 dB and 2.2 dB, respectively, compared to the CPPA.
-This may limit the maximum radar range under STSR mode.
-IV Parameter Sensitivity Analysis
-Figure 13: Pattern performance of the proposed and conventional polarimetric phased arrays under different desired polarization states.
-(a) PSL and XPL under different $\gamma$.
-(b) Pt and Dir under different $\gamma$.
-(c) PSL and XPL under different $\eta$.
-(d) Pt and Dir under different $\eta$.
+10 presents the elevation plane cuts of the CoP and XP patterns for the PCRPA dual-beam array from Fig.
+9, comparing them with the dual-polarized beams from the CPPA and a randomly coded control group.
+The corresponding performance metrics are compared in TABLE IV.
+The control group was configured such that its subarray coding vectors, $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, were randomly generated while maintaining an equal number of excited elements, with $\boldsymbol{x}_{1}=\boldsymbol{x}_{1h}$ and $\boldsymbol{x}_{2}=\boldsymbol{x}_{2v}$.
+This configuration excludes the geometric constraints, element position optimization, and polarization port allocation strategies in Algorithm 2.
+The results demonstrate that the PCRPA optimized by the proposed method maintains comparable PSL while achieving better XPL over the other.
+The XPL improved from -22.46/-22.93 dB to -41.55/-45.78 dB.
+Furthermore, compared to the random control group, the proposed algorithm yields well-matched CoP patterns for the H- and V-beams, as evidenced by their nearly identical main lobe widths in Fig.
+10.
+However, a directivity loss of approximately 2.50 dB is observed in the PCRPA architecture, which is an inherent trade-off due to the use of only a subset of elements for each beam.
+Figure 10: Normalized 2-D power patterns (elevation plane cut) for dual-polarized beams. (a) CoP patterns. (b) XP patterns.
+TABLE IV: Comparison of key metrics for the dual polarized pattern synthesis
+CPPA PCRPA
+random proposed
+XPL (dB) H -22.46 -22.46 -41.55
+V -22.93 -22.94 -45.78
+PSL (dB) H -12.73 -11.68 -13.31
+V -12.75 -8.91 -13.48
+Dir (dBi) H 28.37 25.84 25.87
+V 28.37 26.05 25.86
+IV-C Parameter Sensitivity Analysis
+Figure 11: Comparison of pattern performance between the conventional and proposed architecture under varying polarization states.
+(a)-(d) show XPL, PSL, Dir, and $P_{t}$ as functions of $\gamma$; (e)-(h) present the same metrics as functions of $\eta$.
 The simulation results have verified the effectiveness of the proposed PCRPA architecture and synthesis methods.
 This section explores the sensitivity of pattern performance to various input parameters, including the polarization state, beam direction, and array size.
-IV-A Polarization State
-The desired polarization state affects the pattern performance by modifying the polarization vector’s decomposition results given by Equation (17), and further changing element numbers $n_{h}$, $n_{v}$, and compensation phase $\beta$.
+IV-C1 Polarization State
+This experiment evaluates the impact of polarization parameters on synthesis performance.
+The polarization states for both the CPPA and PCRPA were synthesized by traversing either $\gamma$ or $\eta$ while keeping the other fixed at $\eta=60^{\circ}$ or $\gamma=30^{\circ}$, respectively, with subsequent pattern metrics collected.
+To further quantify the sensitivity to polarization parameters, a new variable $P_{t}$, defined as the radiated power of the CoP pattern at the beam direction, was introduced.
+The results given by Fig.
+11 reveal several key findings.
+First, the CPPA is largely insensitive to polarization variations, whereas the PCRPA exhibits significant sensitivity.
+Second, the PCRPA shows low sensitivity to the polarization angle $\eta$; although the XPL curve in Fig.
+11e exhibits fluctuations, all values remain below -40 dB, and the irregular variations can reasonably be attributed to differences in active element patterns under mutual coupling.
+Crucially, the PCRPA’s performance metrics—PSL, Dir, and $P_{t}$—are highly sensitive to the polarization angle $\gamma$.
+As shown in Figs.
+11b-11d, when $\gamma$ approaches $0^{\circ}$, $45^{\circ}$, or $90^{\circ}$, the patterns display higher PSL and Dir.
+However, a distinct $P_{t}$ loss of approximately 3 dB occurs near $\gamma=45^{\circ}$, while this loss diminishes to nearly zero as $\gamma$ approaches $0^{\circ}$ or $90^{\circ}$.
+This behavior originates from the synthesis mechanism of the PCRPA: the ratio of H-to-V ports is determined by $\gamma$ given by Equation (19).
+When this ratio approaches 1, or when ports are entirely H or V, the PCRPA’s CoP pattern most closely resembles that of the CPPA, leading to similar PSL and Dir.
+However, the vector superposition of orthogonal radiation fields from a nearly equal number of ports results in energy cancellation during polarization synthesis, which causes the observed $P_{t}$ degradation.
+This warrants careful consideration in low SNR environments where power efficiency is critical for detection accuracy.
+IV-C2 Array Size
+Figure 12: Comparison of pattern performance between the conventional and proposed arrays across different array sizes.
+The legend distinguishes the beam types: ’1’ for single-polarized and ’2’ for dual-polarized beams.
+(a) XPL.
+(b) PSL.
+(c) Directivity.
+Figure 13: Comparison of pattern performance between the conventional and proposed arrays across different beam directions.
+The legend distinguishes the beam types: ’1’ for single-polarized and ’2’ for dual-polarized beams.
+(a) XPL.
+(b) PSL.
+(c) Directivity.
+This experiment investigates the impact of array size on synthesis performance.
+With the beam direction and polarization fixed, arrays of sizes $8\times 8$, $16\times 16$, $24\times 24$, and $32\times 32$ were evaluated.
+The AEPs of these arrays served as the input for both CPPA and PCRPA in arbitrary-polarized and dual-polarized pattern synthesis, with subsequent pattern metrics calculated.
+The results given by Fig.
+12 reveals several key trends.
+First, as shown in Fig.
+12a, the XPL of the PCRPA decreases rapidly as the array size increases from $8\times 8$ to $24\times 24$, plateauing after falling below -50 dB.
+Second, Fig.
+12b indicates that the PSL for dual-beam synthesis in the PCRPA improves significantly when scaling from $8\times 8$ to $16\times 16$, then stabilizes, whereas its single-beam PSL declines gradually throughout.
+Furthermore, the directivity of both CPPA and PCRPA increases with array size, following similar growth trends.
+These findings indicate that the XPL and dual-beam PSL of the PCRPA are sensitive to the array size.
+The sensitivity of the dual-beam PSL shares a similar principle with thinned array synthesis, primarily governed by the dimensionality of the optimization problem.
+The sensitivity of the XPL, however, is inherent to its suppression mechanism of the PCRPA.
+The polarization angle $\hat{\gamma}$ in the PCRPA is a discrete, quantized value when mutual coupling is neglected, which can be expressed by:
+$\hat{\gamma}=\arctan(\frac{n_{v}}{n_{h}})$ (30)
+A comparison of the $\hat{\gamma}$ quantization curves for the $8\times 8$ and $32\times 32$ arrays in Fig.
+14 clearly shows that a larger array results in a finer quantization step.
+This enhances the accuracy of approximating the desired polarization state, thereby improving polarization purity and yielding a lower XPL.
+Figure 14: The quantization curves of $\hat{\gamma}$ concerning different array size.
+IV-C3 Beam Direction
+This experiment evaluates the impact of beam direction.
+With the azimuth angle fixed at $\phi=30^{\circ}$, the elevation angle $\theta$ was swept from $30^{\circ}$ to $150^{\circ}$ to define the beam direction for both CPPA and PCRPA in arbitrary-polarized and dual-polarized pattern synthesis.
+The corresponding pattern metrics were recorded.
+The results given by Fig.
+13 reveal several key observations.
+First, the CPPA’s XPL, PSL, and directivity all gradually degrade with beam scanning.
+Second, as shown in Figs.
+13b and 13c, the PSL and directivity of the PCRPA follow the same trend as the CPPA.
+Although the proposed algorithm maintains PSL at a lower level than the CPPA, a consistent directivity loss is observed—approximately 0.3 dB for single-beam and 2.5 dB for dual-beam synthesis.
+Finally, Fig.
+13a demonstrates that the PCRPA achieves significant XPL suppression, with the improvement becoming more pronounced at larger scanning angles.
+The single-beam synthesis yields better XPL suppression compared to the dual-beam case; however, the XPL values themselves fluctuate without a clear correlation to the scanning angle.
+The similarity in PSL and directivity trends between the PCRPA and CPPA is attributed to their shared dependence on the element pattern.
+The XPL suppression in single-beam synthesis occurs because all array DoFs are allocated to synthesizing a single polarization state, resulting in finer equivalent polarization port quantization.
+This principle aligns with the array size sensitivity analysis discussed previously.
+V Experiment and Discussion
+V-A Experimental Results
+To verify the effectiveness of the proposed architecture of the PCRPA and synthesis methods, an experiment is carried out using a standard commercial dual-polarized microstrip patch antenna, shown in Fig.
+15.
+The antenna operates at a frequency of 9.3 GHz, with a bandwidth of 200 MHz.
+The array size of $8\times 8$, and the element spacing is 16.1 mm.
+A planar near-field measurement was conducted for the antenna.
+During the test, the polarized ports were connected via a switch matrix to a vector network analyzer through independent channels, as shown in Fig.
+15b.
+For all these ports, the 3-D AEPs were acquired via probe scanning and near-field to far-field transformation as measured results, which were then used as the input in the simulation platform.
+By controlling the 0-1 coding and phase weighting corresponding to each data set, the physical functions of the RF switches and phase shifters given in Fig.
+2 were emulated, thereby validating the principle of the proposed architecture and methodology.
+Figure 15: A standard commercial X-band $8\times 8$-unit dual-polarized microstrip patch array antenna test in an anechoic chamber. (a) Front view. (b) Back view.
+Figure 16: Measured CoP and XP patterns (elevation cut) under the conventional and proposed architecture for single polarized pattern synthesis.
 Fig.
-13 presents the proposed PCRPA and the CPPA pattern performance concerning different desired polarization states.
-The results are obtained concerning the synthesis of arbitrarily polarized patterns.
-The polarization state parameters are adjusted by varying either $\gamma$ (with $\eta$ fixed at $60^{\circ}$) or $\eta$ (with $\gamma$ fixed at $30^{\circ}$).
-First, compared to the curves shown in Fig.
-13c and 13d, the PCRPA curves in Fig.
-13a and 13b exhibit more pronounced fluctuations, indicating a stronger sensitivity to $\gamma$ than $\eta$.
-Then, considering the impact of $\gamma$, the XPL is consistently suppressed below -40 dB and exhibits little correlation with the polarization state.
-In contrast, the PSL, directivity, and power curves show a more regular trend that changes with $\gamma$.
-When $\gamma$ approaches $45^{\circ}$, the PSL and directivity become close to the CPPA values, but the power loss increases to a maximum of around 3 dB.
-This is because the number of H- and V-polarized elements $n_{h}$ and $n_{v}$ tend to be equal, resulting in similar array factors to those of the CPPA.
-Meanwhile, a significant portion of power is canceled during the vector composing process.
-These can lead to the following conclusions. For a given array size and beam direction:
-  • The XPL in PCRPA pattern synthesis is almost insensitive to the desired polarization state, while the PSL, directivity, and radiated power are sensitive to the auxiliary polarization angle $\gamma$.
-  • The PCRPA exhibits lower PSL and directivity compared to the CPPA.
-Their performance difference gradually diminishes as $\gamma$ approaches $45^{\circ}$, until the characteristics become nearly comparable.
-  • The PCRPA exhibits a certain level of power loss compared to the CPPA.
-The loss magnitude progressively increases as $\gamma$ approaches $45^{\circ}$, with the maximum loss reaching approximately 3 dB.
-IV-B Beam Direction
-Figure 14: Pattern performance of the proposed and conventional polarimetric phased arrays under different beam directions.
-In the legend, ’1’ represents the single-polarized beam, and ’2’ denotes the dual-polarized beam.
-(a) PSL curves.
-(b) XPL curves.
-(c) Directivity curves.
-Figure 15: Pattern performance of the proposed and conventional polarimetric phased arrays under different array sizes.
-In the legend, ’1’ represents the single-polarized beam, and ’2’ denotes the dual-polarized beam.
-(a) PSL curves.
-(b) XPL curves.
-(c) Directivity curves.
-The beam direction affects the pattern performance by determining the specific basis vector sets selected through Equation (15), as well as the element pattern data and weight vectors used in Equation (21).
+16 compares the CoP and XP patterns of the conventional and proposed architectures for single-polarized pattern synthesis, under the same beam direction $(120^{\circ},15^{\circ})$ and V-polarization setting as used in simulations.
+The measured results align well with the simulation trends: the PCRPA achieves a significant XPL suppression from -18.48 dB (CPPA) to -35.26 dB, an improvement of approximately 16.78 dB.
+While this value is slightly higher than the simulated -40 dB level in Fig.
+12b, it still supports the conclusion that the proposed architecture effectively suppresses XPL.
+Furthermore, the PSL remains nearly identical between the two architectures (-13.40 dB for CPPA vs.
+-14.30 dB for PCRPA), and a limited directivity loss of 0.68 dB is observed in the PCRPA.
+These experimental findings—XPL suppression, maintained PSL, and a minor directivity trade-off—collectively validate the key performance trends predicted by the simulations.
+Figure 17: Measured patterns (elevation cut) under different architectures for dual polarized pattern synthesis. (a) H-polarized beam. (b) V-polarized beam.
 Fig.
-14 shows the performance of both single- and dual-polarized pattern synthesis of the PCRPA and the CPPA concerning different beam directions.
-The azimuth angle $\phi$ is fixed at $15^{\circ}$ while the elevation angle $\theta$ ranges from $30^{\circ}$ to $150^{\circ}$ with a step of $15^{\circ}$.
-The polarization state of the single-polarized pattern is fixed at $(30^{\circ},60^{\circ})$.
-Figs.
-14a and 14c reveal the same degradation trends of PSL and directivity for both architectures concerning single- and dual-polarized cases with increasing scanning angles.
-Additionally, the performance gaps between the PCRPA and the CPPA remain stable throughout the scanning range.
-In Fig.
-14b, the XPL of the single-polarized pattern appears insensitive to scanning angle variations.
-In contrast, the dual-beam XPL demonstrates an increasing trend as the scanning angle $\Delta\theta$ enlarges.
-When $\Delta\theta\leq 30^{\circ}$, the dual-beam XPL is lower than the single-beam by 10-20 dB; however, they converge to comparable values when $\Delta\theta>30^{\circ}$.
-This owes to the different XPL suppression mechanisms: the dual-beam employs the calibration matrix with continuously adjustable coefficients, while the single-beam relies on controlling the ratio of H- and V-polarized elements, where the element numbers are discrete.
-In addition, beam pointing error at large scanning angles induces a mismatch in the calibration matrix, deteriorating the XPL in dual-beam synthesis.
-These can lead to the following conclusions. For a given array size and polarization state:
-  • All pattern performance of the PCRPA, except the XPL of the single-polarized beam, exhibits sensitivity to beam scanning, with progressive degradation observed as the scanning angle increases.
-  • The gaps in PSL and directivity between the PCRPA and CPPA remain almost stable throughout the scanning range.
-  • The XPL of the PCRPA is always suppressed at no more than -40 dB, where the dual beam pattern synthesis can exhibit excessively low levels at -60 dB near the broadside.
-IV-C Array Size
-The array size affects the pattern performance by determining the specific number of H- and V-polarized elements in Equation (18), and the dimensionality of the optimization model given by Equation (27).
-As shown in Fig.
-15, all radiation pattern parameters exhibit evident improvements with the array size, except the XPL of dual-beam for the PCRPA.
-Notably, the XPL of dual-beam PCRPA consistently maintains -60 dB-level stability and demonstrates insensitivity to the array size.
-This is because its XP suppression performance is determined by the estimation accuracy of the polarization calibration matrix rather than the array size.
-Particularly, both the PSL of dual-beam synthesis and the XPL of single-beam synthesis show remarkable array size dependence in small-to-medium arrays ($\leq 24\times 24$): The performance enhancement in dual-beam synthesis stems from effective dimensionality expansion of the optimization model.
-In contrast, for the single-beam case, the improvement arises from the quantization characteristics of the polarization angle $\gamma$, given by the equation below:
-$\hat{\gamma}=\arctan(\frac{n_{v}}{n_{h}})$ (29)
-Increasing the array size will reduce the quantization step of $\hat{\gamma}$, thereby improving polarization-state synthesis precision and enhancing XPL performance.
-These can lead to the following conclusions. For a given polarization state and beam direction:
-  • All PCRPA pattern performance, except for the dual-beam XPL, exhibits array-size dependence and shows progressive improvement with the array size.
-  • The XPL performance of PCRPA dual-beam synthesis demonstrates negligible dependence on the array size.
-  • In small-to-medium scale arrays, both the single-beam XPL and dual-beam PSL of the PCRPA demonstrate evident sensitivity to the array size, with rapid performance enhancement observed as the size expands.
-This scaling characteristic strongly suggests the proposed PCRPA architecture for large-scale PPA systems.
-V Experimental Results and Discussion
-V-A Experimental Results
-To verify the effectiveness of the proposed architecture of the PCRPA and synthesis methods, an experiment is carried out using an X-band $8\times 8$-unit dual-polarized microstrip patch array antenna, shown in Fig.
-16.
-The operating frequency is 9.30 GHz.
-The measurement results include active element pattern data for all ports, as shown in Fig.
-16b, which means that the whole array can be utilized as a fully digital system.
-Figure 16: An X-band $8\times 8$-unit dual-polarized microstrip patch array antenna test in an anechoic chamber. (a) Front view. (b) Back view.
-Fig.
-17 shows the CoP and XP patterns under the conventional and proposed architecture for single polarized pattern synthesis, with the beam direction at $(115^{\circ},10^{\circ})$ and the polarization state at V polarization.
-Compared with the CPPA, the PCRPA pattern mitigates the XPL from -19.67 dB to -37.75 dB, improving approximately 18 dB.
-Although this value indicates a degraded performance compared to the figure of simulated cases at -40 dB given in Fig.
-15b, it still agrees with the previous conclusion of XPL suppression.
-Besides, the values of PSL for both conventional and proposed architectures remain comparable at -13.40 dB and -13.99 dB, respectively.
-Figure 17: Measured CoP and XP patterns (elevation cut) under the conventional and proposed architecture for single polarized pattern synthesis.
-Figure 18: Measured patterns (elevation cut) under different architectures for dual polarized pattern synthesis. (a) Conventional PPA. (b) Proposed PCRPA.
-Fig.
-18 shows the 2-D power patterns measured under two architectures for dual-polarized pattern synthesis, with the beam direction the same as the previous single-beam case.
-For the CoP patterns, both architectures demonstrate comparable matching performance: CPPA achieves a 3dB beamwidth of $5.79^{\circ}$ for H- and V-polarized beams, while PCRPA attains a narrower $4.98^{\circ}$ in both polarizations.
-The corresponding matching errors measure 5.35dB (CPPA) and 9.26dB (PCRPA).
-Compared to the CPPA, the CoP patterns of PCRPAs show PSL increase at 1.76 dB and directivity loss at 2.65 dB due to fewer channels.
-In addition, the XPL is effectively suppressed by calibration methods from -19.88 dB to -25.20 dB.
-The experiment results agree with the conclusions drawn by simulations but also reveal performance deterioration due to practical factors.
+17 presents the measured 2-D power patterns for dual-polarized synthesis under the conventional and proposed architectures, also with the beam direction set to $(120^{\circ},15^{\circ})$.
+The experimental outcomes confirm the expected trade-offs: the PCRPA effectively suppresses XPL, reducing it from -14.23/-18.48 dB to -25.62/-39.84 dB through element polarization port reconfiguration.
+However, as also indicated in simulations, this improvement is accompanied by a PSL increase of 3.25 dB and a more noticeable directivity loss of 3.66 dB, attributable to the reduction in active channels.
+It is observed that the H-polarized beam exhibits less XPL suppression than its V-polarized counterpart, a phenomenon we attribute to the heightened sensitivity of mutual coupling, fabrication and environmental variations in smaller arrays.
+It is anticipated that this performance would improve with larger array scales, as stated in the sensitivity analysis, where such practical factors exert a relatively smaller influence.
+Overall, these results align with simulated predictions while underscoring the impact of real-world constraints on achievable performance.
 V-B Discussion
-Analysis of both simulation and experimental data shows that the array size is a crucial factor for the performance of the proposed PCRPA architecture.
-Employing PCRPAs in larger arrays with a minimum size of $8\times 8$ is recommended.
+Analysis of both simulation and experimental results shows that the array size is a crucial factor for the performance of the proposed PCRPA architecture.
+Employing PCRPAs in larger arrays is recommended.
 As the array size increases, the proposed PCRPAs exhibit lower sidelobe levels and better control over cross-polarization levels of the patterns.
@@
   3. 3.
-    For the dual polarized pattern synthesis, larger arrays can offer more DOFs in the thinned array synthesis problem, allowing for higher optimization limits on the PSL of the CoP patterns.
+    For the dual polarized pattern synthesis, larger arrays can offer more DoFs in the thinned array synthesis problem, allowing for higher limits on the PSL of the CoP patterns.
 Furthermore, this architecture’s main advantage is its low cost compared to the conventional design.
@@
 Moreover, due to the reduction in the number of T/R channels, there is an inevitable loss in beam power and directivity.
-Future research will focus on these problems.
+Looking forward, future research will address the identified limitations in PSL and XPL performance.
+Additionally, extending the proposed architecture to conformal arrays presents an important research direction.
+In such geometries, position-dependent polarization variations and complex mutual coupling effects due to surface curvature introduce significant challenges.
+Future work will explore modeling strategies and algorithmic adaptations to handle these problems.
+[Uncaptioned image] Yiqing Wang was born in 1999.
+She received the B.S.
+degree in information engineering from the National University of Defense Technology, Changsha, China, in 2021, where she is currently pursuing the Ph.D.
+degree in information and communication engineering.
+Her research interests include antenna arrays and polarimetric phased arrays.
+[Uncaptioned image] Jian Zhou was born in 1996.
+He received the B.S.
+degree in information engineering and the Ph.D.
+degree in information and communication engineering from the National University of Defense Technology (NUDT), Changsha, China, in 2018 and 2024, respectively.
+In 2025, he joined the College of Electronic Science, NUDT, where he is currently an Assistant Researcher.
+His research interests include antenna arrays and polarimetric phased array.
+[Uncaptioned image] Chen Pang was born in 1986.
+He received the B.S.
+and Ph.D.
+degrees in information and communication engineering from the National University of Defense Technology (NUDT), Changsha, China, in 2009 and 2015, respectively.
+From November 2012 to November 2014, he was a visiting Ph.D.
+Student with the Delft University of Technology, Delft, The Netherlands.
+Currently, he is an Associate Professor with the College of Electronic Science and Technology, National University of Defense Technology.
+His research interests include weather radar signal processing, radar polarimetry, active array, and electromagnetic simulation.
+[Uncaptioned image] Wenyang Man was born in 2000.
+He received the B.S.
+degree in Electronic Engineering from North China Electric Power University, Baoding, China, in 2023, and is currently pursuing the Ph.D.
+degree in Electronic Science and Technology at the National University of Defense Technology, Changsha, China.
+His research interests include shared-aperture polarimetric phased array radar.
+[Uncaptioned image] Zixiang Xiong was born in 2000.
+He received the M.S.
+degree in information engineering from the National University of Defense Technology, Changsha, China, in 2025, where he is currently pursuing the Ph.D.
+degree in electronic information.
+His research interests include antenna arrays and polarimetric phased arrays.
+[Uncaptioned image] Ke Meng was born in 1999.
+He received the B.S.
+degree in information engineering from the National University of Defense Technology, Changsha, China, in 2021, where he is currently pursuing the M.S.
+degree in electronic information.
+His research interests include antenna arrays and polarimetric phased arrays.
+[Uncaptioned image] Zhanling Wang was born in 1988.
+He received the B.S.
+degree in electronic engineering from Shenyang University of Technology, Shenyang, China, in 2007, the M.S.
+degree in information and communication engineering from Air Force Engineering University, Xi’an, China, in 2011, and the Ph.D.
+degree from the National University of Defense Technology, Changsha, China, in 2021.
+He is currently a Lecturer.
+His research interests include array antennas design and polarimetric phased arrays.
+[Uncaptioned image] Yongzhen Li was born in 1977.
+He received the B.S.
+and Ph.D.
+degrees in electronic engineering from the National University of Defense Technology, Changsha, China, in 1999 and 2004, respectively.
+Currently, he is a Professor with the National University of Defense Technology.
+His research interests include radar signal processing, radar polarimetry, and target recognition.